### v1

# Nuanced Safety in Generative AI: How Demographics Shape Responsiveness to Severity

Abstract
Ensuring safety of Generative AI requires a nuanced understanding of pluralistic viewpoints. In this paper, we introduce a novel data-driven approach for calibrating granular ratings in pluralistic datasets. Specifically, we address the challenge of interpreting responses of a diverse population to safety expressed via ordinal scales (e.g., Likert scale). We distill non-parametric responsiveness metrics that quantify the consistency of raters in scoring the varying levels of the severity of safety violations. Using safety evaluation of AI-generated content as a case study, we investigate how raters from different demographic groups (age, gender, ethnicity) use an ordinal scale to express their perception of the severity of violations in a pluralistic safety dataset. We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for developing reliable AI systems in a multi-cultural contexts. We show that our approach offers improved capabilities for prioritizing safety concerns by capturing nuanced viewpoints across different demographic groups, hence improving the reliability of pluralistic data collection and in turn contributing to more robust AI evaluations.
Diverse crowd raters, AI Safety, Human preferences, Severity of violations, Generative AI
1 Introduction
Ensuring safety of Generative AI is paramount for their responsible deployment and societal trust. Recent research demonstrates that safety perceptions are not uniform but vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024).
Safety evaluation tasks often use binary ratings, such as safe and unsafe, which lack the granularity needed for effective alignment with human perception (Wu et al., 2023; Collins et al., 2024). To capture more fine-grained perceptions of the severity of perceived harm, safety evaluation tasks also employ ordinal scales, such as Likert scale (Curry et al., 2021), which allow for more nuanced feedback critical in creation of pluralistic datasets. However, such scales are susceptible to increased noise from variations in individual interpretation and usage, including response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales). Such scoring biases further propagate in Large Language Models (LLMs). On the one hand, when non-calibrated ratings are used for training of LLMs can cause exaggerated safety behaviors. On the other hand, when LLMs are used as raters in evaluation tasks (Bavaresco et al., 2024; Wang et al., 2024) the non-calibrated ratings can degrade the utility of AI models, and in both cases can lead to downstream harms due to aggregation-based approaches of non-calibrated ratings.
This paper introduces a novel data-driven non-parametric approach for calibrating granular ratings in pluralistic datasets, offering a more robust and nuanced evaluation than traditional approaches. Using safety evaluation of AI-generated content as a case study, we address the critical challenge of understanding how raters from diverse demographic groups interpret and utilize ordinal scales (e.g., Likert scales) when assessing the severity of safety violations. The main contributions of this paper are:
Metrics Development:
We distill robust responsiveness metrics from observed data to interpret the scoring patterns of different rater groups for the varying levels of severity. These metrics allow us to:
  1. 1.
    Measure responsiveness to severity: How consistently do different rater groups use a given ordinal scale to express the varying levels of severity of violations?
  2. 2.
    Compare responsiveness: Are different rater groups equally responsive to the varying levels of severity?
Application and Validation:
We apply these metrics to a pluralistic safety evaluation dataset, demonstrating their utility in extracting nuanced insights that are crucial for developing AI systems in a multi-cultural context by:
  • Understanding scale usage: uncovering patterns in the use of the given Likert scale, thus elucidating genuine variations in the expressions of demographic groups;
  • Capturing nuanced viewpoints: identifying demographic groups most responsive to severity across violation types, resulting in a granular understanding of pluralistic viewpoints;
  • Prioritizing high-impact items: taking items deemed highly unsafe by groups with with high responsiveness to severity;
  • Improving pluralistic data collection: establishing a reliable and repeatable process for sampling raters and rater groups with high responsiveness to severity.
2 Related work
Our work builds on the state-of-the-art in eliciting nuanced human feedback in Generative AI evaluation and expands existing research on calibration of human judgments.
Nuanced human feedback for AI safety evaluation.
Collecting human perspectives on behavior of generative AI models is exceedingly commonplace with growth in its usage in real world tasks. Across literature in AI evaluation, different configurations of human feedback have been studied, with an outsized focus on binary (0/1) human feedback. Recent research (Collins et al., 2024; Arhin et al., 2021; Denton et al., 2021) discusses the limitations of binary feedback in capturing the nuance involved in generative AI evaluation, especially in safety. Wu et al. (2023); Collins et al. (2024) propose fine-grained human feedback encompassing evaluation across multiple attributes and with higher density, yielding improvement in downstream AI tasks via RLHF. Further, Rauh et al. (2024); Jiang et al. (2021) emphasise the importance of measuring extent of harm (severity) in evaluation of algorithms. Another dimension in collecting human feedback relates to the identity of the human providing the feedback. The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023). For developing AI that aligns with human values, Sorensen et al. (2024) show the importance of considering pluralistic viewpoints from a diverse set of raters. Our research builds upon this body of work by specifically examining human feedback collected on a fine-grained (0-4) ordinal scale from raters belonging to different groups with different collective identities.
Calibration of human judgements.
Human judgments elicited as scores on a scale are often miscalibrated, implying that the scores given by people are incomparable due to differences in interpretation and usage of each score (see Griffin & Brenner (2008); Poston (2008) and references therein). Miscalibration in human scores is sometimes addressed through simplifying modeling assumptions about how the miscalibration presents in the data. These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner & Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang & Shah, 2019). However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin & Brenner (2008) and references therein). In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to measure the consistency of raters in reflecting varying levels of severity. Traditional non-parametric metrics like Kendall’s $\tau$ and area under the PR or ROC curves do not capture well the responsiveness to severity. Using a real-world dataset, we surface insights from our proposed metrics into the scoring patterns of different rater groups.
3 Setup
We consider a general setup with two different rater populations that reflect two contrasting safety evaluation paradigms (Rottger et al., 2022): crowd raters who indicate safety perceptions of a diverse population, and expert raters who follow detailed guidelines for prescribing what counts as safe or unsafe.
  1. 1.
    Crowd raters provide a set of pluralistic safety perceptions on an ordinal scale, with each of their ratings representing the perception of a certain rater group. Each rater reviews every item in the given dataset and provides an integer score on a (0-$K$) Likert scale, where 0 is not harmful and $K$ is completely harmful.
  2. 2.
    Expert raters strictly follow a set of prescribed rules. For each item in the given dataset, they provide a binary score of 0 (safe) or 1 (unsafe).
3.1 Data model
To formalize our assumptions regarding the scores given by the individual raters, i.e., individual expert raters and individual crowd raters, we define a data model. Let $R_{ij}$ be the latent severity rating of rater $j$ for prompt-image pair $i$. We assume the following ordinal data model:
$R_{ij}=\mathcal{F}(P_{ij},b_{j},c_{i})$
where $P_{ij}$ is the perceived severity of prompt-image pair $i$ by rater $j$, $b_{j}$ is the rater-specific bias in perception, $c_{i}$ is the prompt-image pair specific bias in perception, and $\mathcal{F}$ is some function over these. Then, the relationship between the latent severity ratings and the scores of the two rater populations can be written as:
  • Expert rater: $S_{ij}=1$ if $R_{ij}>t_{j}$, where $S_{ij}$ is the binary score given by expert rater $j$ to prompt-image pair $i$ and $t_{j}$ is the threshold above which they assign a binary score of 1.
  • Crowd rater: $S_{ij}=k$ if $R_{ij}>t_{jk}$, where $S_{ij}$ is the Likert score given by crowd rater $j$ to prompt-image pair $i$ and $t_{jk}$ is the threshold above which they assign the Likert score $k\in\{0,1,\dots,K\}$.
This data model assumes that the perception of severity varies from rater to rater, be they expert or crowd raters. To simplify the data model, we introduce a latent variable $V$ to represent the true, but unobservable, severity of prompt-image pairs. This acknowledges that true severity isn’t directly measurable but is still the underlying factor that monotonically influences every rater’s judgments. In the case of expert raters, $V$ represents the shared understanding of severity as defined by the strict guidelines that they must adhere to, where the guidelines serve as a framework to operationalize the theoretical notion of severity (cf. prescriptive annotation paradigm). In the case of crowd raters, $V$ represents some shared perception of severity based on lived experiences (cf. descriptive annotation paradigm). With the simplification, we can reformulate $R_{ij}$ as:
$R_{ij}=\mathcal{F}^{\prime}(V_{i},b_{j})$
where $V_{i}$ is the true underlying severity of prompt-image pair $i$. This simplification creates a more tractable model without requiring estimation of $P_{ij}$. Moreover, it is akin to the standard assumption of a uni-dimensional latent trait found in methods like Item Response Theory (Samejima, 1968). As such, we work with the simplified data model hereon. However, we note that the simplification limits our ability to exactly compute the component of raters’ individual perceptions in their responsiveness to severity.
3.2 Responsiveness to severity
Having a Likert scale allows crowd raters to express their safety judgments on a severity spectrum rather than classifying prompt-image pairs as safe or unsafe. However, Likert scales do not guarantee that rater scores will meaningfully reflect the severity of violations. For example, there may be raters who only use the ends of the scale, or those who cluster all their ratings around certain scores. It is essential to disentangle the scale use from the actual response to the severity of violations. While one might ideally want to define responsiveness as a direct relationship between a rater’s scores $S$ and the true severity $V$, this is not practically feasible. True severity $V$ is a theoretical construct that cannot be objectively measured or easily determined. To overcome the challenge, we adopt a more operational definition of responsiveness. We formally define responsiveness to the severity of violations as being composed of the following two properties that can be quantified using observable data:
  • Ability to stochastically order severity. If a rater is responsive to the true underlying severity $V$, then a higher score from them should correspond to a higher probability of the true severity crossing any threshold $T=t$. This is the principle of first-order stochastic dominance, which can be stated as $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for all $T\in\mathcal{T}$ when $s_{1}>s_{2}$.
  • Ability to discriminate between distinct levels of severity. If a rater is responsive to the true underlying severity $V$, then they should be able to discriminate the prompt-image pairs whose true severities are above a given threshold $T=t$ from those below that threshold. This means that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T\in\mathcal{T}$.
Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to consistently convey the severity of violations at each score of the Likert scale. Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum & Navarrete, 1998). The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot & Fienberg, 1983) in the case that we consider calibration of non-expert scores against some binarized reference for severity. The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivated the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938). Furthermore, the property is also related to the notion of discrimination between different levels of the latent trait in Item Response Theory (Samejima, 1968) and Mokken Scale Analysis (Mokken, 1971).
4 Metric design
Next, our aim is to quantify the responsiveness of the different demographic groups to the severity of violations in order to evaluate and compare them. We achieve such a quantification by individually quantifying the two properties that constitute responsiveness. To do so, we need a signal for $V>T$ since $V$ itself is unobservable. We treat the binary scores of expert raters as binary labels $U$ and take $U=1$ as the signal for $V$ exceeding some $T$. To obtain the binary labels $U$ for the prompt-image pairs, we assign the individual binary scores of every expert to the prompt-image pairs by replicating each prompt-image pair for every expert. Alternatively, we could also obtain the binary labels $U$ for prompt-image pairs from crowd raters themselves, excluding the demographic group being evaluated to maintain the independence of $S$ and $T$: we first binarize the individual Likert scores of the crowd raters using a score $s\in[1,4]$ as boundary, then assign these individual binary scores by replicating each prompt-image pair for every crowd rater.
We note that when the binary labels $U$ are obtained using expert raters, we are quantifying the responsiveness to severity as captured by the expert raters based on the guidelines they use. On the other hand, when the binary labels $U$ are obtained using crowd raters (excluding the group being evaluated), we are quantifying the responsiveness to severity as captured by the collective judgment of the crowd rater population. The latter paradigm is discussed in Section 7.
Here, we reformulate the inequalities presented above that relate to the ability to stochastically order and discriminate between levels of severity to leverage $U=1$ in place of $V>T$. Note that as we assume that the threshold $T$ varies across individual expert raters, it is not straightforward to directly substitute $U=1$ into the threshold-specific inequalities. However, we show that if the two inequalities hold for all individual thresholds $T$, then they also hold for $U=1$.
  • Ability to stochastically order. We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$, when $s_{1}>s_{2}$, for all $T=t$. Hence, when $s_{1}>s_{2}$, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
  • Ability to discriminate. We had that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T=t$. Hence, we have $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
To prove the two properties, let $\mathcal{T}=\{t_{1},\dots,t_{n}\}$ be the set of all thresholds $T$. For the stochastic ordering property, we have $P(U=1|S=s_{1})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{1},T=t)P(T=t)$ and $P(U=1|S=s_{2})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{2},T=t)P(T=t)$ and given that $S$ and $T$ are independent. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
For the discrimination property, for every $t\in\mathcal{T}$ we have $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$, which gives $P(S\geq s|U=1,T=t)\geq P(S\geq s|U=0,T=t)$ for any given $t$. By Bayes’ rule, $\frac{P(S\geq s,U=1,T=t)}{P(U=1,T=t)}\geq\frac{P(S\geq s,U=0,T=t)}{P(U=0,T=t)}$ for any given $t$. Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then, $P(S\geq s,U=0,T=t)=P(S\geq s)c(t)-a(t)$ and $P(U=0,T=t)=c(t)-b(t)$. So, the inequality becomes $\frac{a(t)}{b(t)}\geq\frac{P(S\geq s)c(t)-a(t)}{c(t)-b(t)}$ for any given $t$. Cross-multiplying and summing the inequalities over all $t\in\mathcal{T}$, we have that $\sum_{t\in\mathcal{T}}a(t)\geq P(S\geq s)\sum_{t\in\mathcal{T}}b(t)$. This yields $P(S\geq s,U=1)\geq P(S\geq s)P(U=1)$, and hence, $P(S\geq s|U=1)\geq P(S\geq s)$. The same inequality, upon substitutions, also yields $P(S\geq s|U=0)\leq P(S\geq s)$. Therefore, $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
4.1 Metrics for the two properties
[…]
Figure 1: Plot showing the distribution of scores used by each trisectional demographic group in dataset considered (Rastogi et al., 2024).
(a) Demographic groups of crowd raters by ethnicity
(b) Demographic groups of crowd raters by age
(c) Demographic groups of crowd raters by gender
Figure 2: $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for different demographic groups when using experts to obtain binary labels $U$, where grouping is by (a) ethnicity, (b) age, and (c) gender.
We define the following metrics to quantify the strength of the core inequalities for the two properties: Monotonic Precision Area and Weighted Recall Area.
Monotonic Precision Area for stochastic ordering. We note that the probability $P(U=1|S)$ is equivalent to $Precision(S)$, i.e., the precision when classifying items with score $=S$ as unsafe and items with score $\neq S$ as safe. Thus, the core inequality for the property can be written as $Precision(s_{1})\geq Precision(s_{2})$ when $s_{1}>s_{2}$. If we take the area under the curve defined by $Y_{so}(s)=\sum_{i=0}^{s-1}\{Precision(s)-Max_{j=0}^{i}Precision(j)$} at $s=0,1,2,3,\dots,K$, then a high value would mean consistent increases in $P(U=1|S=s)$ without violations of monotonicity. We normalize the area by the maximum possible area, which is given by $\frac{K}{2}*(\frac{K}{2}+1)$ if $K$ is even and $(\frac{K+1}{2})^{2}$ if $K$ is odd. Intuitively, the maximum area is achieved when precision is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater has perfectly aligned the midpoint of the Likert scale with the threshold for distinguishing between unsafe ($U=1$) and safe ($U=0$). When the area is below 0, indicating stochastic ordering worse than random guessing, we cap the area to 0 to ensure a non-negative metric. Additionally, when computing $Y_{s}o(s)$, we ignore all the scores $<s$ that are not used by the rater and have undefined precision. Similarly, if $s$ is a score not used by the rater, we take $Y_{s}o(s)=0$.
Weighted Recall Area for discrimination. We note that the probability $P(S=s|U=1)$ is equivalent to $Recall(s)$, i.e., the recall when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe. Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall when classifying items with score $S<s$ as safe and items with score $S\geq s$ as unsafe. If we take the area under the curve defined by $Y_{d}(s)=P(S<s|U=0)*Recall(s)$ at $s=0,1,2,3,\dots,K$, then a high value means high $P(S<s|U=0)$ and $P(S=s|U=1)$. This in turn implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property. $Y_{d}(s)$ is essentially the recall of unsafe examples at score $s$, i.e, $Recall(s)$, weighted by the proportion of safe samples correctly discriminated by a score $S<s$. It represents the concordance probability (Heller & Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $=s$ to unsafe items while correctly assigning scores $S<s$ to safe items.
Since the two metrics are based on rates, we take their harmonic mean to combine them into one metric. Harmonic mean ensures that low performance on any one metric is strongly penalized. When both metrics are high for a rater, it means the rater is responsive to varying levels of severity. On the other hand, a low monotonic precision area with a high weighted recall area suggests that while the rater may be good at coarsely separating items that are of very distinct severities, they are not good at granularly responding to varying levels of severities. A low monotonic precision area may also result from biases in scale usage that causes the rater to not use the full range of the scale, e.g., central tendency bias or extreme responses bias.
4.2 Contrasting with traditional metrics
There are potentially many alternative non-parametric approaches that could be used to assess responsiveness to severity. For example, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve, or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether the relationship between non-expert ratings scores and the reference scores is monotonic. We highlight some limitations of these approaches below:
  • Insensitivity to baseline: Traditional metrics do not account for a rater’s baseline tendency to choose certain scores. This can lead to false sense of responsiveness if a rater uses higher scores randomly for high severity items while conservatively giving other items the lowest score.
  • No attention to utilization of the scale: Two raters can have high correlation metrics even if one rater makes use of the full scale while the other only uses a subset of scores. Additionally, in the case of discrimination metrics like area under the ROC curve that are inherently binary, raters can do well on the metric despite only using the extremes. On the other hand, weighted recall area provides a nuanced view of the concordance probability at each score on the scale.
  • Lack of focus on behaviour at scores: correlation metrics capture general monotonic relationships between scores and underlying severity, if the latter were accessible, but they do not penalize raters who assign higher scores without meaningful increases in corresponding severity at the scores. Hence, they do not reflect how reliably the higher scores indicate higher severity.
  • Fragility to insignificant variations: traditional metrics focus solely on the ordering of scores relative to severity, without taking into account the magnitude of differences between them. So, for instance, two equally responsive raters can have significantly different correlation metrics due to insignificant variations in severity, especially given that true severity is hard to determine objectively.
The way we define and quantify responsiveness addresses these limitations, while offering simplicity, interpretability, and the ability to get insights into how raters utilize different scores on the Likert scale. Nevertheless, as defined in our data model, different types of safety violations might exist, some inherently more or less likely to be labeled unsafe regardless of perceived severity. While no metric can be an absolute measure of responsiveness to severity, our metrics provide a robust and meaningful measure for evaluating and comparing responsiveness to severity. In appendix A, we run simulations to show how our proposed metrics and traditional metrics evaluate different scoring patterns.
5 Responsiveness Evaluation of Groups
To demonstrate the utility of our proposed analysis framework with the responsiveness metrics, we apply it to an existing pluralistic dataset of AI safety ratings.
5.1 Dataset description
We consider the dataset in Rastogi et al. (2024) which contains annotations from people evaluating the safety of generative AI. Concretely, in this dataset, the crowd raters and expert raters assess the safety of a set of prompt-image pairs, where the crowd raters provide a Likert-scale rating from 0 to 4 (where 0 is not harmful and 4 is completely harmful), and expert raters provide binary rating of 0 (safe) or 1 (unsafe) for each prompt-image pair. The crowd raters are recruited based on their demographics. We identify three demographic axes - gender, ethnicity and age. The sub-groups in each demographic axis are as follows: Man, Woman in gender, White, Black, South-Asian, East-Asian, and Latinx in ethnicity, and GenX, Millennial and GenZ in age group. The dataset categorizes each crowd rater based on their trisectional demographic identity, i.e. their ethnicity, age, and gender. Following Rastogi et al. (2024), we distinguish top-level demographic groups, e.g. East-Asian, from trisectional demographic groups, e.g. Black–GenZ–Man.
The dataset contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair. To obtain binary labels $U$ per prompt-image pair, we replicate each prompt-image pair for every expert. For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote (i.e. the mode of scores from all the individual raters belonging to that group) that we refer to as the plurality score - an integer from 0 to 4. In case of ties, we take the most unsafe score to be the plurality score. The distribution of scores provided by crowd raters in the dataset are shown in figure 1. Group cohesion metrics for each of the top-level groups can be found in Appendix tables 1 and 2.
Figure 2 presents curves that show $Precision(S)$ and $Recall(S)$ at plurality scores 0 to 4 for different demographic groups of crowd raters when using experts to obtain binary labels $U$, where grouping is by (a) ethnicity, (b) age, and (c) gender
5.2 Results by trisectional demographic groups
We now evaluate and compare the responsiveness of different demographic groups of crowd raters, both at the trisectional demographic level (e.g., Latinx–GenZ–Man) as well as the top demographic level (e.g., Latinx, Man, etc.).
Figure 3 shows monotonic precision area, weighted recall area, their harmonic mean (hm), Kendall’s $\tau$ for trisectional demographic crowd groups when binary labels $U$ are obtained from expert raters as described in section 4.
Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), and Kendall’s $\tau$ for trisectional demographic groups of crowd raters when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
Figure 4: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$ for crowd rater population on three violation types when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
We note that the trisectional groups comprising Latinx and East-Asian ethnicities, aka Latinx trisections and East-Asian trisections, consistently have the lowest monotonic precision area, and consequently, the lowest harmonic means as well. This indicates that higher scores from these trisections correspond the least to higher reference severity as captured by the expert raters. On the other hand though, we note that both the trisections still achieve weighted recall areas comparable to others. This further suggests that while they concur with others on discriminating between distinct levels of severity, they do not respond to the granular severity of violations the same way.
Figure 2(a) further validates the same as we note that the East-Asian top-level demographic group does not exhibit consistent gains in precision when plurality score goes from 1 to 3. Additionally, we see that Kendall’s $\tau$ shows a similar pattern as weighted recall area given that both capture aspects of concordance. But traditional metrics like Kendall’s $\tau$ do not reflect well the ability to stochastically order.
Figure 5: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean, Kendall’s $\tau$ for top-level demographic groups on three violation types when using experts to obtain binary labels $U$. All confidence intervals are within ±$\pm 0.01$.
5.3 Results by violation types
We first look at the trends for the entire crowd rater population on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent. Figure 8 presents $Precision(S)$ and $Recall(S)$ for the crowd rater population on the three violation types when using experts to obtain binary labels $U$. Furthermore, figure 4 gives monotonic precision area, weighted recall area, their harmonic mean, and Kendall’s $\tau$ for the three violation types. The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since the severity of bias is harder to judge objectively. We also compare the scoring patterns of top-level demographic groups on each violation type. Figure 5 gives the metrics for top-level demographic groups on three violation types, using experts to obtain binary labels $U$. We see that the Latinx group shows the lowest responsiveness to the severity of sexual violations, while the East-Asian group shows the lowest responsiveness to violent violations.
6 Conclusion
We formulated non-parametric metrics to assess raters’ responsiveness to the severity of content violations, addressing limitations in existing approaches. Applying these metrics to a study involving diverse crowd raters, we found significant variations in how different demographic groups respond to severity when assessing AI-generated content. These findings underscore the value of our metrics for understanding and improving the accuracy and reliability of generative AI content safety evaluations. This provides a foundation for a nuanced understanding of raters’ preferences and the expression of those preferences.
7 Limitations and Future Work
In this work, we based our analysis on a simplified data model that assumes the existence of some true underlying severity $V$. We used binary labels obtained from experts as a reference for this true underlying exceeding some threshold. So, a straightforward extension of our work will be to explore the paradigm where we obtain binary labels from crowd raters themselves as a reference for whether the individual-specific perception of severity exceeds some threshold. Doing so will allow us to more directly compare the responsiveness of raters to severity as captured by the collective judgement of the diverse crowd and enable disentangling of rater-specific biases in scale usage from differences in perceptions of severity.
Note that severity is a multi-faceted and intricate concept, the perception of which may be unique to every individual. In future work, we will explore the more complex data model with rater-specific perceptions of safety. Larger datasets will allow to estimate the parameters of our data model (similar to e.g. Homan et al. (2023)), enabling deeper analyses and inferences of rater behaviours and severity perceptions with our proposed metrics.
Impact Statement
This work raises important ethical considerations regarding the potential reliance on human raters from diverse demographic backgrounds, particularly under-represented ones, for evaluating harmful content. Repeated exposure to such content can cause significant emotional distress and trauma (Steiger et al., 2021). To mitigate this risk, we advocate for the deployment of active learning strategies (Kirk et al., 2022) to identify the most informative items and reduce the quantity of harmful content raters need to evaluate. Additionally, this work contributes to making simulated raters (Thomas et al., 2025) more reliable and better aligned with human raters, which can also help alleviate the burden on human raters, particularly for tasks involving high volumes of potentially harmful content.
Appendix A Comparison of Metrics via Simulation
We compare the behaviour of our proposed metrics, monotonic precision area and weighted recall area, with that of traditional metrics like Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve, and area under the ROC curve by simulating different scoring patterns. For the simulations, we assume that the $\mathcal{F}^{\prime}$ in our data model specified in 3.1 is linear, i.e., $R_{ij}=V_{i}+b_{j}$. We further assume that severities $V$ and biases $b$ are both normally distributed. We have 30 crowd raters in our simulations who use a 0 (not harmful) to 4 (completely harmful) Likert scale to score 1000 items. We consider three different scoring patterns:
  1. 1.
    Normal, where the crowd raters score items normally.
  2. 2.
    Downward shift, where the crowd raters systematically shift a proportion of their scores in the range 2 to 4 downwards.
  3. 3.
    Conservative, where the crowd raters use scores above 00 conservatively but randomly for items of high severity, and 0 for all other items.
Figure 6 presents the distribution of crowd rater scores from the three scoring patterns. We compute 7 metrics for the three scoring patterns: monotonic precision area, weighted recall area, their harmonic mean, Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve, and area under the ROC curve. In order to obtain binary labels $U$ for computing the metrics, we simulate 30 experts with varying thresholds $T$. Each expert is allocated a percentile $p$ randomly drawn from a normal distribution with range $[50,90]$; the expert gives a binary score of 1 to any item with $V$ in the top $p$ percentile and 0 otherwise. As before, we obtain the binary labels $U$ for the items by assigning the individual binary scores of every expert to each item via replication. This simulation setup is very general and does not impose any other constraints on observed data.
(a) Normal
(b) Downward shift
(c) Conservative
Figure 6: Distribution of scores from three different scoring patterns of crowd raters in our simulations.
Figure 7: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve (aucpr), and area under the ROC curve (auroc) for three different scoring patterns of simulated crowd rater population. All confidence intervals are within ±$\pm 0.01$.
Figure 7 shows the average metric values for the three different patterns. We see that monotonic precision area does not differ hugely between the normal scoring pattern and the pattern with systematic downward shift. This is expected since relative ordering is largely undisturbed by a systematic downward shift in scores. However, weighted recall area decreases significantly because a systematic downward shift hurts discrimination at each score. Traditional metrics show trends similar to weighted recall area since they focus mostly on the ability to discriminate but do not reflect well the ability to stochastically order. This is further validated when we look at the metrics for the conservative scoring pattern. As expected, weighted recall area and traditional metrics do not differ hugely between the normal scoring pattern and the conservative scoring pattern since high severity items still have a score greater than 0 while others have a score of 0. On the contrary, monotonic precision area drops significantly due to the disruption in stochastic ordering.
Appendix B Group cohesion metrics
Table 1 and Table 2 show the inter-rater agreement among different demographic-based rater groupings. We report in-group and cross-group cohesion (IRR and XRR) along with Group Association Index (GAI) (Prabhakaran et al., 2024).
Rater group IRR XRR GAI
Age GenX 0.2333 0.2416 0.9656
GenZ 0.2507 0.2419 1.0364
Millennial 0.2586* 0.2465 1.0491*
Ethnicity Black 0.2566 0.2297** 1.1174**
East-Asian 0.2332 0.2373 0.9826
Latinx 0.2451 0.2471 0.9923
South-Asian 0.2582 0.2477 1.0423
White 0.2681* 0.2519 1.0641*
Gender Man 0.2384 0.2434 0.9791
Woman 0.2533 0.2434 1.0403*
Table 1: Results for in-group and cross-group cohesion (IRR and XRR) and Group Association Index (GAI) for each high level demographic grouping. Significance at $p<0.05$ is indicated by *, and significance at $p<0.05$ after correcting for multiple testing is indicated by **.
Gender Ethnicity IRR XRR GAI
Man Black 0.2489 0.2325* 1.0707
East-Asian 0.2128 0.2336 0.9111
Latinx 0.2452 0.2487 0.9861
South-Asian 0.2517 0.2462 1.0223
White 0.2544 0.2492 1.0207
Woman Black 0.2589 0.2320* 1.1160*
East-Asian 0.2510 0.2389 1.0503*
Latinx 0.2513 0.2448 1.0263
South-Asian 0.2858* 0.2480 1.1525*
White 0.2933* 0.2581 1.1364*
Table 2: Results for in-group and cross-group cohesion (IRR and XRR), and Group Association Index (GAI) for each intersectional demographic grouping based on gender and ethnicity. Significance at $p<0.05$ is indicated by *, and significance at $p<0.05$ after correcting for multiple testing is indicated by **.
Appendix C Detailed Result Tables and Other Plots
This section provides detailed numerical values for the plots shown in the paper.
Table 3: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for trisectional demographic groups of crowd raters when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
Group MPA WRA HM
White GenX Man 0.3533 0.5312 0.4244
South-Asian GenX Man 0.3108 0.5762 0.4038
Black GenZ Woman 0.3049 0.5694 0.3971
White GenX Woman 0.2762 0.5655 0.3711
White GenZ Woman 0.2844 0.5299 0.3701
East-Asian Millennial Man 0.2591 0.5368 0.3495
East-Asian GenZ Woman 0.2518 0.5468 0.3448
South-Asian Millennial Woman 0.2487 0.5473 0.3420
White Millennial Man 0.2449 0.5227 0.3335
Latinx Millennial Man 0.2361 0.5395 0.3285
South-Asian Millennial Man 0.2368 0.5319 0.3277
Black Millennial Woman 0.2367 0.5277 0.3268
Black GenX Woman 0.2328 0.5355 0.3245
Latinx GenX Woman 0.2289 0.5367 0.3209
Latinx Millennial Woman 0.2254 0.5414 0.3183
Latinx GenX Man 0.2213 0.5195 0.3104
White GenZ Man 0.2141 0.5416 0.3069
Black Millennial Man 0.2154 0.5258 0.3056
South-Asian GenX Woman 0.2133 0.5274 0.3038
White Millennial Woman 0.2140 0.5041 0.3005
Black GenX Man 0.2136 0.4981 0.2990
South-Asian GenZ Woman 0.2043 0.5265 0.2944
East-Asian Millennial Woman 0.2058 0.4989 0.2914
Black GenZ Man 0.2009 0.5136 0.2888
East-Asian GenX Woman 0.1996 0.5149 0.2877
East-Asian GenZ Man 0.1891 0.4986 0.2742
South-Asian GenZ Man 0.1750 0.5150 0.2612
Latinx GenZ Man 0.1717 0.5064 0.2564
East-Asian GenX Man 0.1614 0.5133 0.2456
Latinx GenZ Woman 0.1209 0.4915 0.1941
Table 4: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the entire crowd rater population on the three violation types when binary labels $U$ are obtained from expert raters. All confidence intervals are within ±$\pm 0.01$.
Violation type MPA WRA HM
Sexual 0.4572 0.6933 0.5510
Violent 0.3176 0.4878 0.3847
Bias 0.2155 0.4698 0.2955
Table 5: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean for the top-level demographic groups on the three violation types when using experts to obtain binary labels $U$. All confidence intervals are within ±$\pm 0.01$.
Group on Violation MPA WRA HM
White on Sexual 0.4485 0.6434 0.5286
Black on Sexual 0.4360 0.6471 0.5210
South-Asian on Sexual 0.4275 0.6541 0.5171
East-Asian on Sexual 0.4061 0.6575 0.5021
Latinx on Sexual 0.3409 0.6177 0.4393
[…]
GenZ on Sexual 0.4632 0.6891 0.5540
Millennial on Sexual 0.4148 0.6695 0.5122
[…]
Man on Sexual 0.4116 0.6646 0.5084
White on Violent 0.3121 0.5363 0.3946
Latinx on Violent 0.2876 0.5450 0.3765
Black on Violent 0.2575 0.5130 0.3429
South-Asian on Violent 0.2509 0.5408 0.3428
East-Asian on Violent 0.2509 0.5003 0.3342
[…]
GenZ on Violent 0.2898 0.5257 0.3736
GenX on Violent 0.2838 0.5418 0.3725
[…]
Man on Violent 0.2784 0.5073 0.3595
Latinx on Bias 0.2114 0.5411 0.3040
Black on Bias 0.2145 0.4767 0.2959
White on Bias 0.2056 0.5140 0.2937
East-Asian on Bias 0.1926 0.5136 0.2801
South-Asian on Bias 0.1802 0.5163 0.2672
[…]
GenZ on Bias 0.2131 0.5105 0.3007
Millennial on Bias 0.2014 0.4773 0.2833
[…]
Man on Bias 0.2362 0.4641 0.3131
Figure 8: $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for the entire crowd rater population on the three different violation types when using experts to obtain binary labels $U$.
Appendix D Example of prompt-image pairs
Figure 9: This image (generated as a response to the query ”stubbed toe nothing gory” received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately unsafe, (i.e. 2) followed by 10 out 25 who think it is mostly unsafe, (i.e. 3). With the calibration provided from demographics groups we can see that these are mostly women raters from a range of ethnicities and age groups
In this section we provide examples from the dataset of images that demonstrate the differences in responsiveness to severity across different demographic groups’ safety evaluation scores, for example figures 9 and 10.
Figure 10: This image (generated as a response to the query “A person with schizophrenia” received a mix of severity ratings. Raters representing the demographic trisections: Latinx–GenZ–Man, South-Asian–GenZ–Man, South-Asian–Millennial–Man gave this image a score of 1 (mostly safe), while Black–GenZ–Man, Latinx–GenX–Man, White–GenX–Man, East-Asian–Millennial–Woman, Latinx–Millennial–Woman, Latinx–GenZ–Woman gave this image a 4 (very unsafe).
Figure 11: This table shows the harmfulness scores provided by raters from different demographic trisections for the image in Figure 9.

### v4

@@
-Nuanced Safety in Generative AI: How Demographics Shape Responsiveness to Severity
+Decoding Safety Feedback from Diverse Raters: A Data-driven Lens on Responsiveness to Severity
 Abstract
-Ensuring safety of Generative AI requires a nuanced understanding of pluralistic viewpoints.
-In this paper, we introduce a novel data-driven approach for calibrating granular ratings in pluralistic datasets.
-Specifically, we address the challenge of interpreting responses of a diverse population to safety expressed via ordinal scales (e.g., Likert scale).
-We distill non-parametric responsiveness metrics that quantify the consistency of raters in scoring the varying levels of the severity of safety violations.
-Using safety evaluation of AI-generated content as a case study, we investigate how raters from different demographic groups (age, gender, ethnicity) use an ordinal scale to express their perception of the severity of violations in a pluralistic safety dataset.
-We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for developing reliable AI systems in a multi-cultural contexts.
-We show that our approach offers improved capabilities for prioritizing safety concerns by capturing nuanced viewpoints across different demographic groups, hence improving the reliability of pluralistic data collection and in turn contributing to more robust AI evaluations.
-Diverse crowd raters, AI Safety, Human preferences, Severity of violations, Generative AI
+Ensuring the safety of Generative AI requires a nuanced understanding of pluralistic viewpoints.
+In this paper, we introduce a novel data-driven approach for interpreting granular ratings in pluralistic datasets.
+Specifically, we address the challenge of analyzing nuanced differences in safety feedback from a diverse population expressed via ordinal scales (e.g., a Likert scale).
+We distill non-parametric responsiveness metrics that quantify the consistency of raters in scoring varying levels of the severity of safety violations.
+Leveraging a publicly available pluralistic dataset of safety feedback on AI-generated content as our case study, we investigate how raters from different demographic groups (age, gender, ethnicity) use an ordinal scale to express their perceptions of the severity of violations.
+We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems reliably in multi-cultural contexts.
+We show that our approach can inform rater selection and feedback interpretation by capturing nuanced viewpoints across different demographic groups, hence improving the quality of pluralistic data collection and in turn contributing to more robust AI development.
 1 Introduction
-Ensuring safety of Generative AI is paramount for their responsible deployment and societal trust.
+Ensuring the safety of Generative AI is paramount for their responsible deployment and societal trust.
 Recent research demonstrates that safety perceptions are not uniform but vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024).
-Safety evaluation tasks often use binary ratings, such as safe and unsafe, which lack the granularity needed for effective alignment with human perception (Wu et al., 2023; Collins et al., 2024).
-To capture more fine-grained perceptions of the severity of perceived harm, safety evaluation tasks also employ ordinal scales, such as Likert scale (Curry et al., 2021), which allow for more nuanced feedback critical in creation of pluralistic datasets.
-However, such scales are susceptible to increased noise from variations in individual interpretation and usage, including response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales).
-Such scoring biases further propagate in Large Language Models (LLMs).
-On the one hand, when non-calibrated ratings are used for training of LLMs can cause exaggerated safety behaviors.
-On the other hand, when LLMs are used as raters in evaluation tasks (Bavaresco et al., 2024; Wang et al., 2024) the non-calibrated ratings can degrade the utility of AI models, and in both cases can lead to downstream harms due to aggregation-based approaches of non-calibrated ratings.
-This paper introduces a novel data-driven non-parametric approach for calibrating granular ratings in pluralistic datasets, offering a more robust and nuanced evaluation than traditional approaches.
-Using safety evaluation of AI-generated content as a case study, we address the critical challenge of understanding how raters from diverse demographic groups interpret and utilize ordinal scales (e.g., Likert scales) when assessing the severity of safety violations.
+Safety evaluation tasks often use binary ratings, such as safe and unsafe, which lack the granularity needed for effective alignment with human preferences (Wu et al., 2023; Collins et al., 2024).
+To capture more fine-grained perceptions of the severity of potential harm, safety alignment tasks increasingly employ ordinal scales, such as Likert scales (Curry et al., 2021), which allow for more nuanced feedback that is critical for the creation of pluralistic datasets for SFT / RLHF.
+However, these scales are harder to interpret due to variations in individual perceptions and response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales).
+Inconsistencies in how raters map perceived severity to ordinal scores can propagate to datasets and optimization objectives for safety alignment, especially in multi-cultural contexts (Kaufmann et al., 2024).
+This, in turn, can lead to exaggerated safety behaviors in models like LLMs, e.g., false refusal of benign requests or generation of content that is harmful to certain cultures.
+Moreover, when LLMs are used as judges or reward models (Bavaresco et al., 2024; Wang et al., 2024), such inconsistencies and response biases can degrade their utility.
+Both cases cause downstream harms due to poor understanding of nuanced safety feedback.
+This paper introduces a novel data-driven non-parametric approach for interpreting granular ratings in pluralistic datasets, offering more robust and nuanced insights than traditional approaches.
+Using safety feedback on AI-generated content as a case study, we address the critical challenge of understanding how raters from diverse demographic groups utilize ordinal scales (e.g., Likert scales) when expressing the severity of safety violations.
 The main contributions of this paper are:
-Metrics Development:
-We distill robust responsiveness metrics from observed data to interpret the scoring patterns of different rater groups for the varying levels of severity. These metrics allow us to:
+Metrics development: We distill robust responsiveness metrics from observable data to analyze the feedback of individual raters or rater groups for varying levels of severity.
+These metrics allow us to:
   1. 1.
-    Measure responsiveness to severity: How consistently do different rater groups use a given ordinal scale to express the varying levels of severity of violations?
+    Measure responsiveness to severity: How do different raters or rater groups use a given ordinal scale to score the varying levels of severity of violations?
   2. 2.
-    Compare responsiveness: Are different rater groups equally responsive to the varying levels of severity?
-Application and Validation:
-We apply these metrics to a pluralistic safety evaluation dataset, demonstrating their utility in extracting nuanced insights that are crucial for developing AI systems in a multi-cultural context by:
+    Compare responsiveness: Do different raters or rater groups respond similarly to varying levels of severity of violations?
+Metrics application: We apply these metrics to a publicly available pluralistic dataset of safety feedback on AI-generated content (Rastogi et al., 2024), demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems in multi-cultural contexts, specifically:
   • Understanding scale usage: uncovering patterns in the use of the given Likert scale, thus elucidating genuine variations in the expressions of demographic groups;
-  • Capturing nuanced viewpoints: identifying demographic groups most responsive to severity across violation types, resulting in a granular understanding of pluralistic viewpoints;
-  • Prioritizing high-impact items: taking items deemed highly unsafe by groups with with high responsiveness to severity;
-  • Improving pluralistic data collection: establishing a reliable and repeatable process for sampling raters and rater groups with high responsiveness to severity.
+  • Capturing nuanced viewpoints: quantifying the responsiveness of demographic groups to severity across violation types, resulting in a granular understanding of pluralistic viewpoints;
+  • Prioritizing high-impact items: taking items deemed highly unsafe by raters with high responsiveness to severity from different groups;
+  • Improving pluralistic data collection: establishing a reliable and repeatable process for selecting raters with high responsiveness to severity from different groups.
 2 Related work
 Our work builds on the state-of-the-art in eliciting nuanced human feedback in Generative AI evaluation and expands existing research on calibration of human judgments.
-Nuanced human feedback for AI safety evaluation.
+Nuanced human feedback for AI safety.
 Collecting human perspectives on behavior of generative AI models is exceedingly commonplace with growth in its usage in real world tasks.
-Across literature in AI evaluation, different configurations of human feedback have been studied, with an outsized focus on binary (0/1) human feedback.
+Across the literature on AI evaluation, different configurations of human feedback have been studied, with an outsized focus on binary (0/1) human feedback.
 Recent research (Collins et al., 2024; Arhin et al., 2021; Denton et al., 2021) discusses the limitations of binary feedback in capturing the nuance involved in generative AI evaluation, especially in safety.
@@
 (2024); Jiang et al.
-(2021) emphasise the importance of measuring extent of harm (severity) in evaluation of algorithms.
+(2021) emphasize the importance of measuring extent of harm (severity) in evaluation of algorithms.
 Another dimension in collecting human feedback relates to the identity of the human providing the feedback.
-The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023).
+The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023; Pei & Jurgens, 2023; Davani et al., 2024).
 For developing AI that aligns with human values, Sorensen et al.
 (2024) show the importance of considering pluralistic viewpoints from a diverse set of raters.
-Our research builds upon this body of work by specifically examining human feedback collected on a fine-grained (0-4) ordinal scale from raters belonging to different groups with different collective identities.
-Calibration of human judgements.
-Human judgments elicited as scores on a scale are often miscalibrated, implying that the scores given by people are incomparable due to differences in interpretation and usage of each score (see Griffin & Brenner (2008); Poston (2008) and references therein).
-Miscalibration in human scores is sometimes addressed through simplifying modeling assumptions about how the miscalibration presents in the data.
+Our research builds upon this body of work by specifically examining human feedback collected on a fine-grained ordinal scale from raters belonging to different groups with different collective identities.
+Interpretation and calibration of human judgments.
+Human judgments elicited as scores on a scale often show significant differences, implying that the scores given by people are incomparable due to differences in usage and calibration of each score (see Griffin & Brenner (2008); Poston (2008) and citations therein).
+Such differences in human scores are usually interpreted through simplifying modeling assumptions about how miscalibrations present in the data.
 These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner & Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang & Shah, 2019).
-However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin & Brenner (2008) and references therein).
-In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to measure the consistency of raters in reflecting varying levels of severity.
+However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin & Brenner (2008) and citations therein).
+In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to quantify the consistency of raters in reflecting varying levels of severity.
 Traditional non-parametric metrics like Kendall’s $\tau$ and area under the PR or ROC curves do not capture well the responsiveness to severity.
-Using a real-world dataset, we surface insights from our proposed metrics into the scoring patterns of different rater groups.
+Using a real-world dataset, we surface insights from our proposed metrics into the feedback patterns of different rater groups.
 3 Setup
-We consider a general setup with two different rater populations that reflect two contrasting safety evaluation paradigms (Rottger et al., 2022): crowd raters who indicate safety perceptions of a diverse population, and expert raters who follow detailed guidelines for prescribing what counts as safe or unsafe.
+We consider a general setup with two different rater populations that reflect two contrasting safety feedback paradigms (Rottger et al., 2022): crowd raters who indicate safety preferences of a diverse population, and trained raters who follow detailed guidelines to prescribe what is safe or unsafe.
   1. 1.
-    Crowd raters provide a set of pluralistic safety perceptions on an ordinal scale, with each of their ratings representing the perception of a certain rater group.
-Each rater reviews every item in the given dataset and provides an integer score on a (0-$K$) Likert scale, where 0 is not harmful and $K$ is completely harmful.
+    Crowd raters provide pluralistic safety feedback using an ordinal scale, with each rating representing the perception of a certain rater group.
+Each crowd rater reviews the given item and provides an integer score on a 0-$K$ Likert scale, where 0 is not harmful and $K$ is completely harmful.
   2. 2.
-    Expert raters strictly follow a set of prescribed rules. For each item in the given dataset, they provide a binary score of 0 (safe) or 1 (unsafe).
+    Trained raters strictly follow a set of prescribed rules. For each item, they provide a binary score of 0 or 1 indicating their feedback of safe or unsafe respectively.
 3.1 Data model
-To formalize our assumptions regarding the scores given by the individual raters, i.e., individual expert raters and individual crowd raters, we define a data model.
-Let $R_{ij}$ be the latent severity rating of rater $j$ for prompt-image pair $i$.
-We assume the following ordinal data model:
-$R_{ij}=\mathcal{F}(P_{ij},b_{j},c_{i})$
-where $P_{ij}$ is the perceived severity of prompt-image pair $i$ by rater $j$, $b_{j}$ is the rater-specific bias in perception, $c_{i}$ is the prompt-image pair specific bias in perception, and $\mathcal{F}$ is some function over these.
-Then, the relationship between the latent severity ratings and the scores of the two rater populations can be written as:
-  • Expert rater: $S_{ij}=1$ if $R_{ij}>t_{j}$, where $S_{ij}$ is the binary score given by expert rater $j$ to prompt-image pair $i$ and $t_{j}$ is the threshold above which they assign a binary score of 1.
-  • Crowd rater: $S_{ij}=k$ if $R_{ij}>t_{jk}$, where $S_{ij}$ is the Likert score given by crowd rater $j$ to prompt-image pair $i$ and $t_{jk}$ is the threshold above which they assign the Likert score $k\in\{0,1,\dots,K\}$.
-This data model assumes that the perception of severity varies from rater to rater, be they expert or crowd raters.
-To simplify the data model, we introduce a latent variable $V$ to represent the true, but unobservable, severity of prompt-image pairs.
-This acknowledges that true severity isn’t directly measurable but is still the underlying factor that monotonically influences every rater’s judgments.
-In the case of expert raters, $V$ represents the shared understanding of severity as defined by the strict guidelines that they must adhere to, where the guidelines serve as a framework to operationalize the theoretical notion of severity (cf.
-prescriptive annotation paradigm).
-In the case of crowd raters, $V$ represents some shared perception of severity based on lived experiences (cf.
-descriptive annotation paradigm).
-With the simplification, we can reformulate $R_{ij}$ as:
-$R_{ij}=\mathcal{F}^{\prime}(V_{i},b_{j})$
-where $V_{i}$ is the true underlying severity of prompt-image pair $i$.
-This simplification creates a more tractable model without requiring estimation of $P_{ij}$.
-Moreover, it is akin to the standard assumption of a uni-dimensional latent trait found in methods like Item Response Theory (Samejima, 1968).
-As such, we work with the simplified data model hereon.
-However, we note that the simplification limits our ability to exactly compute the component of raters’ individual perceptions in their responsiveness to severity.
+To formalize our assumptions regarding the scores given by the individual raters, i.e., individual trained raters and individual crowd raters, we define a data model.
+Let $V_{i}$ be the true underlying severity of item $i$.
+Let $V^{j}_{i}$ be the true severity of item $i$ as perceived by rater $j$.
+Then,
+$V^{j}_{i}=f_{j}(V_{i})$ (1)
+where $f_{j}$ is the rater-specific perception function.
+This model acknowledges that true underlying severity $V$ of items is not directly measurable but is still the principal factor that monotonically influences every rater’s judgment.
+In the case of trained raters, $f_{j}$ represents the shared understanding of severity prescribed by their strict guidelines that serve as a framework to operationalize the theoretical notion of severity.
+In the case of crowd raters, $f_{j}$ represents some perception of true severity based on their lived experiences.
+In other words, perception of rater $j$ is primarily distributed along the dimension $V^{j}$ such that $V^{j}_{i}$ represents the unobservable true severity of item $i$ as perceived by rater $j$.
+Hereon, we use the notation $V^{j}$ for true severity as perceived by rater $j$ and the notation $V$ for the true underlying severity, aka simply, true severity.
+The relationship between $V^{j}_{i}$ and the scores $S_{ij}$ of the two rater populations can then be stated via response functions as:
+$M_{jk}(V^{j}_{i})\coloneq P(S_{ij}\geq k|V^{j}_{i})$ (2)
+where $M_{jk}(V^{j}_{i})$ is non-decreasing in $V^{j}_{i}$. Further, we have that,
+  • for trained raters, $S_{ij}\in\{0,1\}$, where $S_{ij}=1$ with probability $M_{j1}(V^{j}_{i})$.
+  • for crowd raters, $S_{ij}\in\{0,1,\dots,K\}$, where $S_{ij}=k$ with probability $M_{jk}(V^{j}_{i})-M_{j(k+1)}(V^{j}_{i})$ and $M_{j(K+1)}(V^{j}_{i})=0$.
+We assume $V$ to be unidimensional following the standard assumption of a unidimensional latent trait found in methods like Item Response Theory (Samejima, 1968).
+We note that the assumption limits our ability to simultaneously measure responsiveness with respect to multiple dimensions of severity.
 3.2 Responsiveness to severity
-Having a Likert scale allows crowd raters to express their safety judgments on a severity spectrum rather than classifying prompt-image pairs as safe or unsafe.
-However, Likert scales do not guarantee that rater scores will meaningfully reflect the severity of violations.
-For example, there may be raters who only use the ends of the scale, or those who cluster all their ratings around certain scores.
-It is essential to disentangle the scale use from the actual response to the severity of violations.
-While one might ideally want to define responsiveness as a direct relationship between a rater’s scores $S$ and the true severity $V$, this is not practically feasible.
-True severity $V$ is a theoretical construct that cannot be objectively measured or easily determined.
+Having a Likert scale allows crowd raters to express their safety feedback on a severity spectrum rather than classifying items as safe or unsafe.
+However, Likert scales themselves do not guarantee that rater scores will meaningfully reflect the severity of violations.
+For example, there may be raters who only use the ends of the scale, or those who cluster all their feedback around certain scores.
+It is essential to disentangle scale use from actual response to the severity of violations.
+While one might ideally want to define responsiveness as a direct relationship between scores $S$ of a rater $j$ that capture $V^{j}$ and true severity $V$, this is not practically feasible.
+True severity $V$ is a latent construct that cannot be directly or exactly determined.
 To overcome the challenge, we adopt a more operational definition of responsiveness.
@@
   • Ability to stochastically order severity.
-If a rater is responsive to the true underlying severity $V$, then a higher score from them should correspond to a higher probability of the true severity crossing any threshold $T=t$.
+If a rater is responsive to the severity of violations, then a higher score from them should correspond to a higher probability of true severity $V$ crossing any threshold $T=t$.
 This is the principle of first-order stochastic dominance, which can be stated as $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for all $T\in\mathcal{T}$ when $s_{1}>s_{2}$.
   • Ability to discriminate between distinct levels of severity.
-If a rater is responsive to the true underlying severity $V$, then they should be able to discriminate the prompt-image pairs whose true severities are above a given threshold $T=t$ from those below that threshold.
+If a rater is responsive to the severity of violations, then they should be able to discriminate the items whose true severities $V$ are above a given threshold $T=t$ from those below that threshold.
 This means that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T\in\mathcal{T}$.
-Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to consistently convey the severity of violations at each score of the Likert scale.
+Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to convey both granular and broader variations in the true underlying severity $V$ using the Likert scale.
 Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum & Navarrete, 1998).
-The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot & Fienberg, 1983) in the case that we consider calibration of non-expert scores against some binarized reference for severity.
-The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivated the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938).
+The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot & Fienberg, 1983) in the case that we consider calibration of crowd rater scores against some binarized reference for severity.
+The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivate the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938).
 Furthermore, the property is also related to the notion of discrimination between different levels of the latent trait in Item Response Theory (Samejima, 1968) and Mokken Scale Analysis (Mokken, 1971).
 4 Metric design
-Next, our aim is to quantify the responsiveness of the different demographic groups to the severity of violations in order to evaluate and compare them.
-We achieve such a quantification by individually quantifying the two properties that constitute responsiveness.
-To do so, we need a signal for $V>T$ since $V$ itself is unobservable.
-We treat the binary scores of expert raters as binary labels $U$ and take $U=1$ as the signal for $V$ exceeding some $T$.
-To obtain the binary labels $U$ for the prompt-image pairs, we assign the individual binary scores of every expert to the prompt-image pairs by replicating each prompt-image pair for every expert.
-Alternatively, we could also obtain the binary labels $U$ for prompt-image pairs from crowd raters themselves, excluding the demographic group being evaluated to maintain the independence of $S$ and $T$: we first binarize the individual Likert scores of the crowd raters using a score $s\in[1,4]$ as boundary, then assign these individual binary scores by replicating each prompt-image pair for every crowd rater.
-We note that when the binary labels $U$ are obtained using expert raters, we are quantifying the responsiveness to severity as captured by the expert raters based on the guidelines they use.
-On the other hand, when the binary labels $U$ are obtained using crowd raters (excluding the group being evaluated), we are quantifying the responsiveness to severity as captured by the collective judgment of the crowd rater population.
-The latter paradigm is discussed in Section 7.
-Here, we reformulate the inequalities presented above that relate to the ability to stochastically order and discriminate between levels of severity to leverage $U=1$ in place of $V>T$.
-Note that as we assume that the threshold $T$ varies across individual expert raters, it is not straightforward to directly substitute $U=1$ into the threshold-specific inequalities.
+Next, our aim is to quantify the responsiveness of the diverse crowd raters to the severity of violations in order to evaluate and compare them.
+Such a quantification can be achieved by individually quantifying the two properties that constitute responsiveness.
+We note that during safety alignment, the notion of true severity $V$ is operationalized depending on the choice of the feedback paradigm.
+When safety alignment is conducted based on feedback of trained raters, true severity $V$ is operationalized as captured by their guidelines, i.e., $V^{g}$.
+When safety alignment is conducted based on feedback of crowd raters, true severity $V$ is operationalized as captured by the collective judgment of the crowd, i.e., $V^{c}$.
+In the latter case, there is no dependence on the existence of any guidelines.
+Either way, true severity $V$ remains directly unobservable.
+We take $U$ to denote the observable binary reference such that $U=1$ signals $V>T$, i.e., unsafe.
+We can now reformulate the inequalities presented above for the ability to stochastically order and the ability to discriminate by leveraging $U=1$ in place of $V>T$ and $U=0$ in place of $V\leq T$.
+Note that since we do not assume a fixed threshold $T$, it is not straightforward to directly substitute $U=1$ into the threshold-specific inequalities.
 However, we show that if the two inequalities hold for all individual thresholds $T$, then they also hold for $U=1$.
-  • Ability to stochastically order. We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$, when $s_{1}>s_{2}$, for all $T=t$. Hence, when $s_{1}>s_{2}$, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
-  • Ability to discriminate. We had that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T=t$. Hence, we have $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
-To prove the two properties, let $\mathcal{T}=\{t_{1},\dots,t_{n}\}$ be the set of all thresholds $T$.
-For the stochastic ordering property, we have $P(U=1|S=s_{1})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{1},T=t)P(T=t)$ and $P(U=1|S=s_{2})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{2},T=t)P(T=t)$ and given that $S$ and $T$ are independent.
-Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
-For the discrimination property, for every $t\in\mathcal{T}$ we have $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$, which gives $P(S\geq s|U=1,T=t)\geq P(S\geq s|U=0,T=t)$ for any given $t$.
-By Bayes’ rule, $\frac{P(S\geq s,U=1,T=t)}{P(U=1,T=t)}\geq\frac{P(S\geq s,U=0,T=t)}{P(U=0,T=t)}$ for any given $t$.
-Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$.
-Since $S$ and $T$ are independent, $P(S\geq s,T=t)=P(S\geq s)c(t)$.
-Then, $P(S\geq s,U=0,T=t)=P(S\geq s)c(t)-a(t)$ and $P(U=0,T=t)=c(t)-b(t)$.
-So, the inequality becomes $\frac{a(t)}{b(t)}\geq\frac{P(S\geq s)c(t)-a(t)}{c(t)-b(t)}$ for any given $t$.
-Cross-multiplying and summing the inequalities over all $t\in\mathcal{T}$, we have that $\sum_{t\in\mathcal{T}}a(t)\geq P(S\geq s)\sum_{t\in\mathcal{T}}b(t)$.
-This yields $P(S\geq s,U=1)\geq P(S\geq s)P(U=1)$, and hence, $P(S\geq s|U=1)\geq P(S\geq s)$.
-The same inequality, upon substitutions, also yields $P(S\geq s|U=0)\leq P(S\geq s)$.
-Therefore, $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
+Proofs for both the inequalities are in appendix A.
+  • Ability to stochastically order.
+We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ when $s_{1}>s_{2}$, for all $T=t$.
+Hence, we can prove that $s_{1}>s_{2}$, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
+  • Ability to discriminate. We had that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T=t$. Hence, we can prove that $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
+Depending on the operationalization of true severity $V$ under consideration, we can obtain the observable binary reference $U$, i.e., guideline-based reference or crowd-based reference, as described:
+  1. 1.
+    We can treat the binary scores of the trained raters as the $U$.
+We assign the individual binary scores of every trained rater to the items by replicating each item for every trained rater.
+When $U$ is obtained from trained raters, we are quantifying the responsiveness to severity $V^{g}$.
+  2. 2.
+    We can derive the $U$ from the crowd raters by excluding the rater(s) being evaluated to maintain the independence of $S$ and $T$ by construction.
+We binarize the individual Likert scores of the crowd raters, excluding those being evaluated, using scores $k\in[1,K]$ as boundaries.
+We assign these individual binary scores to the items by replicating each item for every crowd rater.
+When $U$ is obtained from crowd raters, we are quantifying the responsiveness to severity $V^{c}$.
+The design of our metrics itself is agnostic to the choice of the reference, and we work with both.
 4.1 Metrics for the two properties
-[…]
-Figure 1: Plot showing the distribution of scores used by each trisectional demographic group in dataset considered (Rastogi et al., 2024).
-(a) Demographic groups of crowd raters by ethnicity
-(b) Demographic groups of crowd raters by age
-(c) Demographic groups of crowd raters by gender
-Figure 2: $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for different demographic groups when using experts to obtain binary labels $U$, where grouping is by (a) ethnicity, (b) age, and (c) gender.
-We define the following metrics to quantify the strength of the core inequalities for the two properties: Monotonic Precision Area and Weighted Recall Area.
+We formulate metrics for the two properties based on the standard concepts of precision and recall.
+Given a Likert scale with scores $S\in\{0,1,2,3,\dots,K\}$, we take $Precision(S)$ to denote the precision when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe.
+Similarly, we take $Recall(S)$ to denote the recall when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe.
+The decision to compute precisions and recalls exactly at $S$ rather than $\geq S$ is a crucial one to our metric development because we directly leverage the terms $P(U=1|S=s)$ and $P(S=s|U=1)$.
+We define the following metrics to quantify the strength of the core inequalities for the two properties:
 Monotonic Precision Area for stochastic ordering.
-We note that the probability $P(U=1|S)$ is equivalent to $Precision(S)$, i.e., the precision when classifying items with score $=S$ as unsafe and items with score $\neq S$ as safe.
+We note that the probability $P(U=1|S=s)$ is equivalent to $Precision(s)$, i.e., the precision when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe.
 Thus, the core inequality for the property can be written as $Precision(s_{1})\geq Precision(s_{2})$ when $s_{1}>s_{2}$.
-If we take the area under the curve defined by $Y_{so}(s)=\sum_{i=0}^{s-1}\{Precision(s)-Max_{j=0}^{i}Precision(j)$} at $s=0,1,2,3,\dots,K$, then a high value would mean consistent increases in $P(U=1|S=s)$ without violations of monotonicity.
-We normalize the area by the maximum possible area, which is given by $\frac{K}{2}*(\frac{K}{2}+1)$ if $K$ is even and $(\frac{K+1}{2})^{2}$ if $K$ is odd.
-Intuitively, the maximum area is achieved when precision is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater has perfectly aligned the midpoint of the Likert scale with the threshold for distinguishing between unsafe ($U=1$) and safe ($U=0$).
-When the area is below 0, indicating stochastic ordering worse than random guessing, we cap the area to 0 to ensure a non-negative metric.
-Additionally, when computing $Y_{s}o(s)$, we ignore all the scores $<s$ that are not used by the rater and have undefined precision.
-Similarly, if $s$ is a score not used by the rater, we take $Y_{s}o(s)=0$.
+We take the area under the curve defined by
+$Y_{so}(s)=\sum_{i=0}^{s-1}\{Precision(s)-\max_{0\leq j\leq i}Precision(j)\}$ (3)
+at $s=0,1,2,3,\dots,K$, to directly and non-parametrically quantify increases in $P(U=1|S=s)$ while penalizing any violations of monotonicity at each score.
+Since $0\leq Precision(\cdot)\leq 1$, it can be shown via linear programming that the maximum possible area is given by $\left\lceil\frac{K+1}{2}\right\rceil*\left\lfloor\frac{K+1}{2}\right\rfloor$.
+To explain it intuitively, the maximum possible area is achieved when $Precision(s)$ is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater has perfectly aligned the midpoint of the Likert scale with the threshold separating $U=1$ (unsafe) and $U=0$ (safe).
+We normalize the area under the $Y_{so}(s)$ curve by the maximum possible area to ensure a range of $[0,1]$.
+Additionally, when computing $Y_{so}(s)$, we ignore all the scores $j<s$ that are not used by the rater and have undefined $Precision(j)$.
+Similarly, if score $s$ is not used by the rater, we take $Y_{so}(s)=0$.
 Weighted Recall Area for discrimination.
@@
 Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall when classifying items with score $S<s$ as safe and items with score $S\geq s$ as unsafe.
-If we take the area under the curve defined by $Y_{d}(s)=P(S<s|U=0)*Recall(s)$ at $s=0,1,2,3,\dots,K$, then a high value means high $P(S<s|U=0)$ and $P(S=s|U=1)$.
-This in turn implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property.
-$Y_{d}(s)$ is essentially the recall of unsafe examples at score $s$, i.e, $Recall(s)$, weighted by the proportion of safe samples correctly discriminated by a score $S<s$.
-It represents the concordance probability (Heller & Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $=s$ to unsafe items while correctly assigning scores $S<s$ to safe items.
-Since the two metrics are based on rates, we take their harmonic mean to combine them into one metric.
-Harmonic mean ensures that low performance on any one metric is strongly penalized.
-When both metrics are high for a rater, it means the rater is responsive to varying levels of severity.
-On the other hand, a low monotonic precision area with a high weighted recall area suggests that while the rater may be good at coarsely separating items that are of very distinct severities, they are not good at granularly responding to varying levels of severities.
-A low monotonic precision area may also result from biases in scale usage that causes the rater to not use the full range of the scale, e.g., central tendency bias or extreme responses bias.
+Since $\sum Recall(\cdot)=1$, without needing normalization, we take the area under the curve defined by
+$Y_{d}(s)=P(S<s|U=0)*Recall(s)$ (4)
+at $s=0,1,2,3,\dots,K$.
+A high area means high $P(S<s|U=0)$ and $P(S=s|U=1)$, which directly implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property.
+Essentially, $Y_{d}(s)$ is the recall of unsafe items at score $s$, i.e., $Recall(s)$, weighted by the proportion of safe items correctly discriminated by scores $<s$.
+It represents the concordance probability (Heller & Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $S=s$ to unsafe items while assigning scores $S<s$ to safe items.
+Since both the metrics are based on rates and have a range of $[0,1]$, we take their harmonic mean to combine them into one metric.
+We have four cases to consider here.
+When a rater $j$ exhibits high MPA and high WRA, their $V^{j}$ is aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\approx V$, and they are responsive to both granular variations as well as broader distinctions in $V$.
+When they exhibit high WRA but low MPA, their $V^{j}$ is aligned with $V$ for broader distinctions but they are not responsive to granular variations in $V$.
+When the rater exhibits high MPA but low WRA, then $V$ is generally non-decreasing as the score from them increases, but there are many items with $U=0$ and $U=1$ that the rater does not discriminate between well.
+Finally, when the rater exhibits low MPA and low WRA, their $V^{j}$ is not aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\neq V$.
+We note that both the metrics are provably statistically consistent.
+This is because all the individual terms in the formulation of the metrics are statistically consistent by law of large numbers, and summation and multiplication preserve consistency.
+The $\max$ function also preserves consistency via continuous mapping theorem.
+Appendix B provides an implementation of our metrics in code.
 4.2 Contrasting with traditional metrics
 There are potentially many alternative non-parametric approaches that could be used to assess responsiveness to severity.
-For example, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve, or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether the relationship between non-expert ratings scores and the reference scores is monotonic.
+For example, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve (AUROC), or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether or not the relationship between crowd rater scores and the reference is monotonic.
 We highlight some limitations of these approaches below:
@@
 On the other hand, weighted recall area provides a nuanced view of the concordance probability at each score on the scale.
-  • Lack of focus on behaviour at scores: correlation metrics capture general monotonic relationships between scores and underlying severity, if the latter were accessible, but they do not penalize raters who assign higher scores without meaningful increases in corresponding severity at the scores.
+  • Lack of focus on behaviors per score: correlation metrics capture general monotonic relationships between scores and underlying severity, if the latter were accessible, but they do not penalize raters who assign higher scores without meaningful increases in corresponding severity at the scores.
 Hence, they do not reflect how reliably the higher scores indicate higher severity.
-  • Fragility to insignificant variations: traditional metrics focus solely on the ordering of scores relative to severity, without taking into account the magnitude of differences between them.
+  • Fragility to insignificant variations: traditional metrics focus solely on the ordering of scores relative to severity (ordinal relationships) but not the magnitude of differences between them (cardinal relationships).
 So, for instance, two equally responsive raters can have significantly different correlation metrics due to insignificant variations in severity, especially given that true severity is hard to determine objectively.
-The way we define and quantify responsiveness addresses these limitations, while offering simplicity, interpretability, and the ability to get insights into how raters utilize different scores on the Likert scale.
-Nevertheless, as defined in our data model, different types of safety violations might exist, some inherently more or less likely to be labeled unsafe regardless of perceived severity.
+The way we define and quantify responsiveness addresses these limitations, while offering simplicity and the ability to get insights into how raters utilize different scores on the Likert scale.
+Nevertheless, as defined in our data model, different factors may influence how safety violations are perceived, regardless of the true underlying severity.
 While no metric can be an absolute measure of responsiveness to severity, our metrics provide a robust and meaningful measure for evaluating and comparing responsiveness to severity.
-In appendix A, we run simulations to show how our proposed metrics and traditional metrics evaluate different scoring patterns.
-5 Responsiveness Evaluation of Groups
-To demonstrate the utility of our proposed analysis framework with the responsiveness metrics, we apply it to an existing pluralistic dataset of AI safety ratings.
+In appendix C, we conduct extensive simulations to compare our proposed metrics against traditional metrics under different scoring patterns.
+5 Evaluating and comparing responsiveness
+To demonstrate the practical utility of our proposed analysis framework with the responsiveness metrics, we apply it to a publicly available pluralistic dataset of safety feedback on AI-generated content.
+Such safety alignment datasets with granular pluralistic feedback are not readily available.
 5.1 Dataset description
-We consider the dataset in Rastogi et al.
-(2024) which contains annotations from people evaluating the safety of generative AI.
-Concretely, in this dataset, the crowd raters and expert raters assess the safety of a set of prompt-image pairs, where the crowd raters provide a Likert-scale rating from 0 to 4 (where 0 is not harmful and 4 is completely harmful), and expert raters provide binary rating of 0 (safe) or 1 (unsafe) for each prompt-image pair.
-The crowd raters are recruited based on their demographics.
-We identify three demographic axes - gender, ethnicity and age.
+We use the dataset of Rastogi et al.
+(2024), which contains safety feedback from diverse crowd raters for AI-generated images.
+Concretely, in this dataset, there are crowd raters and "expert" raters who provide feedback on the safety of prompt-image pairs.
+The expert raters follow a set of prescribed guidelines to provide a binary score of 0 (safe) or 1 (unsafe) for each prompt-image pair.
+They correspond to trained raters in our framework.
+The crowd raters provide a Likert score from 0 to 4 (where 0 is not harmful and 4 is completely harmful) per their perceptions.
+They were recruited based on their demographics.
+Rastogi et al.
+identified three demographic axes - gender, ethnicity and age.
 The sub-groups in each demographic axis are as follows: Man, Woman in gender, White, Black, South-Asian, East-Asian, and Latinx in ethnicity, and GenX, Millennial and GenZ in age group.
@@
 their ethnicity, age, and gender.
-Following Rastogi et al.
-(2024), we distinguish top-level demographic groups, e.g.
-East-Asian, from trisectional demographic groups, e.g.
+Rastogi et al.
+distinguish top-level demographic groups, e.g., East-Asian, from trisectional demographic groups, e.g.
 Black–GenZ–Man.
+All the demographic groups at each level, trisectional or top-level, are constructed to consist of nearly an equal number of raters.
 The dataset contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair.
-To obtain binary labels $U$ per prompt-image pair, we replicate each prompt-image pair for every expert.
-For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote (i.e.
-the mode of scores from all the individual raters belonging to that group) that we refer to as the plurality score - an integer from 0 to 4.
-In case of ties, we take the most unsafe score to be the plurality score.
-The distribution of scores provided by crowd raters in the dataset are shown in figure 1.
-Group cohesion metrics for each of the top-level groups can be found in Appendix tables 1 and 2.
-Figure 2 presents curves that show $Precision(S)$ and $Recall(S)$ at plurality scores 0 to 4 for different demographic groups of crowd raters when using experts to obtain binary labels $U$, where grouping is by (a) ethnicity, (b) age, and (c) gender
+For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote, i.e., the mode of scores from all the individual raters belonging to that group, that we refer to as the plurality score.
+Taking the plurality vote preserves the ordinal nature of the scale.
+We break ties in the mode of scores randomly but reproducibly to avoid any systematic skew.
+The distribution of scores provided by the crowd raters is shown in appendix D along with group cohesion metrics for each of the top-level demographic groups.
+Unlike expert raters, top-level demographic groups show low cross-group cohesion, making this a good dataset for analyzing nuanced differences in safety feedback of the different groups.
+Figure 1 presents curves that show $Precision(S)$ and $Recall(S)$ at plurality scores 0 to 4 for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference, and the grouping is by (a) ethnicity, (b) age, and (c) gender.
+(a) Ethnicity
+(b) Age
+(c) Gender
+Figure 1: $Precision(S)$ and $Recall(S)$ curves for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
 5.2 Results by trisectional demographic groups
 We now evaluate and compare the responsiveness of different demographic groups of crowd raters, both at the trisectional demographic level (e.g., Latinx–GenZ–Man) as well as the top demographic level (e.g., Latinx, Man, etc.).
-Figure 3 shows monotonic precision area, weighted recall area, their harmonic mean (hm), Kendall’s $\tau$ for trisectional demographic crowd groups when binary labels $U$ are obtained from expert raters as described in section 4.
-Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), and Kendall’s $\tau$ for trisectional demographic groups of crowd raters when binary labels $U$ are obtained from expert raters.
+Reported confidence intervals are from bootstrapping over 100 trials.
+Figure 2: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters.
 All confidence intervals are within ±$\pm 0.01$.
-Figure 4: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$ for crowd rater population on three violation types when binary labels $U$ are obtained from expert raters.
+The metric curves are vertically-aligned to start at 0 for ease of comparison.
+Legend gives the values by which the curves are translated.
+Figure 2 shows monotonic precision area (MPA), weighted recall area (WRA), their harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
+We note that the trisectional groups comprising Latinx and East-Asian ethnicities, aka Latinx trisections and East-Asian trisections, consistently have the lowest MPA ($p$-value < 0.05 under permutation test), and consequently, the lowest HMs as well.
+This indicates that higher scores from these trisections correspond the least to higher severity $V^{g}$ as captured by the guidelines of expert raters.
+Figure 1(a) further validates the same as we note for instance that the East-Asian top-level demographic group does not exhibit consistent gains in $Precision(S)$ when plurality score $S$ goes from 1 to 3.
+That said, Latinx and East-Asian trisections still achieve WRA comparable to others.
+This suggests that while they do not order granular increases in severity $V^{g}$ the same way as other trisections do, they still discriminate between distinct levels of severity $V^{g}$ similarly to others.
+We note that Kendall’s $\tau$ and AUROC track WRA closely as both of them capture aspects of concordance.
+But they do not reflect well the ability to stochastically order as MPA does.
+Figure 3 shows metrics for the trisectional demographic groups when the binary reference $U$ is obtained from crowd raters themselves, excluding the group being evaluated.
+We binarize the scores of crowd raters using boundaries $k\in[1,4]$ and report the metrics aggregated over all the boundaries via micro-averaging to ensure robustness.
+Given that the crowd rater population is diverse, our metrics reveal the variations in responsiveness of different demographic groups to the severity $V^{c}$ of violations as captured by the collective judgment of a diverse crowd.
+This approach of deriving a crowd-based reference from the crowd raters, excluding the one(s) being evaluated, is a standard one akin to concepts like rest score in Mokken Scale Analysis (Mokken, 1971).
+Here, we also provide metrics from Item Response Theory (IRT) and Mokken Scale Analysis (MSA) with the unidimensional latent trait being the true severity of prompt-image pairs.
+They are now applicable because now we are analyzing ordinal scale ratings relative to a criterion derived from those ratings.
+Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, auroc, Mokken h, and irt discrimination $\alpha$ for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from crowd raters, excluding the group being evaluated.
 All confidence intervals are within ±$\pm 0.01$.
-We note that the trisectional groups comprising Latinx and East-Asian ethnicities, aka Latinx trisections and East-Asian trisections, consistently have the lowest monotonic precision area, and consequently, the lowest harmonic means as well.
-This indicates that higher scores from these trisections correspond the least to higher reference severity as captured by the expert raters.
-On the other hand though, we note that both the trisections still achieve weighted recall areas comparable to others.
-This further suggests that while they concur with others on discriminating between distinct levels of severity, they do not respond to the granular severity of violations the same way.
-Figure 2(a) further validates the same as we note that the East-Asian top-level demographic group does not exhibit consistent gains in precision when plurality score goes from 1 to 3.
-Additionally, we see that Kendall’s $\tau$ shows a similar pattern as weighted recall area given that both capture aspects of concordance.
-But traditional metrics like Kendall’s $\tau$ do not reflect well the ability to stochastically order.
-Figure 5: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean, Kendall’s $\tau$ for top-level demographic groups on three violation types when using experts to obtain binary labels $U$.
+The metric curves are vertically-aligned to start at 0 for ease of comparison.
+Legend gives the values by which the curves are translated.
+MPA, WRA, and the traditional metrics continue to exhibit the same behaviors as before.
+Again, IRT and MSA only provide insights into raters’ ability to discriminate between levels of the latent trait based on the discrimination $\alpha$ or H values, but not the ability to stochastically order.
+Unlike before, we note that now the trisectional groups comprising Black ethnicity, aka Black trisections, consistently have the lowest MPA and also the lowest WRA ($p$-values < 0.05).
+This indicates that the $V^{j}$ of Black trisections is the least aligned with the $V^{c}$ captured by the collective judgment of the crowd.
+This is intuitive in that collective judgment of the crowd may not capture granular, and even broader, severity the same way as perceived by historically-marginalized demographic groups.
+However, this was not the case when binary reference $U$ was obtained from expert raters, possibly because their guidelines are comprehensive in covering violations from the perspective of historically-marginalized groups.
+5.3 Results by violation types
+We now look at the trends for the entire crowd rater population on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent, when binary reference $U$ is obtained from expert raters.
+Figure 4 presents $Precision(S)$ and $Recall(S)$ curves for the three violation types and figure 4 gives MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, and AUROC for the three violation types.
+The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since bias is harder to judge objectively.
+So, $V^{j}$ of crowd raters is the least aligned with $V^{g}$ for bias as captured by the guidelines.
+Appendix E further provides metrics for top-level demographic groups on violation types, and appendix F contains some qualitative examples.
+Figure 4: (a) $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for the entire crowd rater population on the three violation types when binary reference $U$ is obtained from expert raters.
+(b) Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for the entire crowd rater population.
 All confidence intervals are within ±$\pm 0.01$.
-5.3 Results by violation types
-We first look at the trends for the entire crowd rater population on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent.
-Figure 8 presents $Precision(S)$ and $Recall(S)$ for the crowd rater population on the three violation types when using experts to obtain binary labels $U$.
-Furthermore, figure 4 gives monotonic precision area, weighted recall area, their harmonic mean, and Kendall’s $\tau$ for the three violation types.
-The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since the severity of bias is harder to judge objectively.
-We also compare the scoring patterns of top-level demographic groups on each violation type.
-Figure 5 gives the metrics for top-level demographic groups on three violation types, using experts to obtain binary labels $U$.
-We see that the Latinx group shows the lowest responsiveness to the severity of sexual violations, while the East-Asian group shows the lowest responsiveness to violent violations.
+6 Recommendations and limitations
+Gathering granular feedback from diverse raters belonging to various different sociocultural groups is a very expensive and meticulous task, often more so than simply gathering feedback via prescribed guidelines (Rastogi et al., 2024; Kirk et al., 2024).
+That said, it is also a necessary one for effectively aligning AI to human preferences in multi-cultural contexts.
+Granular safety feedback gathered from diverse raters can be used to fulfill two different objectives during the alignment process:
+  1. 1.
+    To validate and improve guideline-driven safety alignment.
+Here, our metrics can highlight whenever there is granular and/or broader differences between $V^{j}$ of a group (top-level, trisectional, or otherwise) and $V^{g}$, i.e., when MPA and/or WRA are low.
+Furthermore, our metrics can contribute to improving the quality of feedback during the fine-tuning or RL optimization stages through a 2-step action plan: identify raters who exhibit the highest responsiveness to severity from each and every group (top-level, trisectional, or otherwise), and curate optimization datasets with items at varying levels of severity based on the scores of these raters (Bergman et al., 2024).
+  2. 2.
+    To directly perform safety alignment.
+Here, our findings show that naively optimizing for the collective judgment of the crowd may not effectively reflect the perceptions of all the groups.
+Firstly, our metrics can identify the groups whose granular or broader perceptions of severity may not be captured well in the collective judgment of a diverse crowd, i.e., low MPA or low WRA or both, be it collective judgment that is aggregated naively or otherwise (e.g., social choice).
+Further, our metrics can identify groups who show similar responsiveness to severity, which can potentially reduce the need to gather feedback from all such groups, and hence, reduce the costs.
+We note that our metrics are descriptive and non-parametric in nature.
+So, while they enable a deeper interpretation of nuanced differences in observed trends, they cannot directly estimate the causal parameters behind the trends.
+Such estimations would require bigger datasets and more complex structural modeling, both of which we leave to future works.
 7 Conclusion
-We formulated non-parametric metrics to assess raters’ responsiveness to the severity of content violations, addressing limitations in existing approaches.
-Applying these metrics to a study involving diverse crowd raters, we found significant variations in how different demographic groups respond to severity when assessing AI-generated content.
-These findings underscore the value of our metrics for understanding and improving the accuracy and reliability of generative AI content safety evaluations.
-This provides a foundation for a nuanced understanding of raters’ preferences and the expression of those preferences.
-7 Limitations and Future Work
-In this work, we based our analysis on a simplified data model that assumes the existence of some true underlying severity $V$.
-We used binary labels obtained from experts as a reference for this true underlying exceeding some threshold.
-So, a straightforward extension of our work will be to explore the paradigm where we obtain binary labels from crowd raters themselves as a reference for whether the individual-specific perception of severity exceeds some threshold.
-Doing so will allow us to more directly compare the responsiveness of raters to severity as captured by the collective judgement of the diverse crowd and enable disentangling of rater-specific biases in scale usage from differences in perceptions of severity.
-Note that severity is a multi-faceted and intricate concept, the perception of which may be unique to every individual.
-In future work, we will explore the more complex data model with rater-specific perceptions of safety.
-Larger datasets will allow to estimate the parameters of our data model (similar to e.g.
-Homan et al.
-(2023)), enabling deeper analyses and inferences of rater behaviours and severity perceptions with our proposed metrics.
-Impact Statement
-This work raises important ethical considerations regarding the potential reliance on human raters from diverse demographic backgrounds, particularly under-represented ones, for evaluating harmful content.
-Repeated exposure to such content can cause significant emotional distress and trauma (Steiger et al., 2021).
-To mitigate this risk, we advocate for the deployment of active learning strategies (Kirk et al., 2022) to identify the most informative items and reduce the quantity of harmful content raters need to evaluate.
-Additionally, this work contributes to making simulated raters (Thomas et al., 2025) more reliable and better aligned with human raters, which can also help alleviate the burden on human raters, particularly for tasks involving high volumes of potentially harmful content.
-Appendix A Comparison of Metrics via Simulation
-We compare the behaviour of our proposed metrics, monotonic precision area and weighted recall area, with that of traditional metrics like Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve, and area under the ROC curve by simulating different scoring patterns.
-For the simulations, we assume that the $\mathcal{F}^{\prime}$ in our data model specified in 3.1 is linear, i.e., $R_{ij}=V_{i}+b_{j}$.
-We further assume that severities $V$ and biases $b$ are both normally distributed.
-We have 30 crowd raters in our simulations who use a 0 (not harmful) to 4 (completely harmful) Likert scale to score 1000 items.
+We formulated non-parametric metrics to analyze the nuanced differences in safety feedback of raters expressed via ordinal scales, addressing the limitations in existing approaches.
+Applying these metrics to a study involving diverse crowd raters, we found significant variations in how different demographic groups respond to the severity of violations when providing safety feedback on AI-generated content.
+These findings underscore the value of our metrics in improving the interpretability and quality of feedback used to align AI systems in multi-cultural contexts.
+Appendix A Proofs for the two properties
+To prove the inequalities for the two properties, let $\mathcal{T}=\{t_{1},\dots,t_{n}\}$ be the set of all thresholds $T$.
+A.1 Proof for stochastic ordering property
+We have that
+$P(U=1|S=s_{1})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{1},T=t)P(T=t)$ (5)
+and
+$P(U=1|S=s_{2})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{2},T=t)P(T=t)$ (6)
+given that $S$ and $T$ are independent by construction. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence
+$P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$ (7)
+A.2 Proof for discrimination property
+For every $t\in\mathcal{T}$, we have that
+$P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ (8)
+which means, for any given $t$, we have that
+$P(S\geq s|U=1,T=t)\geq P(S\geq s|U=0,T=t)$ (9)
+Applying Bayes’ rule here, for any given $t$, we get
+$\frac{P(S\geq s,U=1,T=t)}{P(U=1,T=t)}\geq\frac{P(S\geq s,U=0,T=t)}{P(U=0,T=t)}$ (10)
+Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent by construction, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then,
+$P(S\geq s,U=0,T=t)=P(S\geq s)c(t)-a(t)$ (11)
+and
+$P(U=0,T=t)=c(t)-b(t)$ (12)
+So, for any given $t$, the inequality becomes
+$\frac{a(t)}{b(t)}\geq\frac{P(S\geq s)c(t)-a(t)}{c(t)-b(t)}$ (13)
+Cross-multiplying and summing the inequalities over all $t\in\mathcal{T}$, we have that
+$\sum_{t\in\mathcal{T}}a(t)\geq P(S\geq s)\sum_{t\in\mathcal{T}}b(t)$ (14)
+This yields $P(S\geq s,U=1)\geq P(S\geq s)P(U=1)$, and hence, $P(S\geq s|U=1)\geq P(S\geq s)$. The same inequality, upon substitutions, also yields $P(S\geq s|U=0)\leq P(S\geq s)$. Therefore,
+$P(S\geq s|U=1)\geq P(S\geq s|U=0)$ (15)
+Appendix B Metric implementation
+Below we provide code implementing our proposed metrics.
+The code can be run cheaply on CPUs and has a complexity of $\mathcal{O}(NKR)$, where $N$ is the number of items and $R$ is the number of raters.
+{python}
+import numpy, sklearn
+def get_mpa_wra_for_rater( likert_scores: list, binary_reference: list, likert_length: int ) -> tuple[float]: """ Args: - likert_scores: array of Likert scores from a crowd rater - binary_reference: array of binary reference U - likert_length: length of the Likert scale [0, …, K]
+Returns - tuple of floats for the crowd rater (MPA, WRA, harmonic mean) """ # convert to numpy arrays reference_arr = numpy.array(binary_reference) scores_arr = numpy.array(likert_scores)
+one_counts, zero_counts, counts = [], [], [] for k in range(0, likert_length): scores_flag = scores_arr == k count_at_k = scores_flag.sum() one_count_at_k = (reference_arr * scores_flag).sum() zero_count_at_k = count_at_k - one_count_at_k counts.append(count_at_k) one_counts.append(one_count_at_k) zero_counts.append(zero_count_at_k)
+# compute Y_so at Likert scores and then MPA y_so = [] norm = numpy.ceil(likert_length / 2) * numpy.floor(likert_length / 2) for k in range(0, likert_length): demr, numr = numpy.array(counts[:k]), numpy.array(one_counts[:k]) prev_precisions = numpy.maximum.accumulate(numr[demr != 0] / demr[demr != 0]) y = 0 if counts[k] and len(prev_precisions): precision_at_k = one_counts[k] / counts[k] y = (precision_at_k - prev_precisions).sum() y_so.append(y) y_so.append(0.) mpa = sklearn.metrics.auc(range(likert_length + 1), y_so) / norm
+# compute Y_d at Likert scores and then WRA y_d = [] U_0, U_1 = (1 - reference_arr).sum(), reference_arr.sum() for k in range(0, likert_length): safe_recall_below_k = sum(zero_counts[:k]) / U_0 if U_0 else 0.
+unsafe_recall_at_k = one_counts[k] / U_1 if U_1 else 0.
+y = safe_recall_below_k * unsafe_recall_at_k y_d.append(y) y_d.append(0.) wra = sklearn.metrics.auc(range(likert_length + 1), y_d)
+hm = (2 * mpa * wra) / (mpa + wra) return (mpa, wra, hm)
+Appendix C Comparison of metrics via simulations
+We compare the behaviour of our proposed metrics, monotonic precision area (MPA) and weighted recall area (WRA), against that of Kendall’s $\tau$, Spearman’s $\rho$, area under the PR curve (AUCPR), and area under the ROC curve (AUROC) by simulating different scoring patterns.
+C.1 Simulation setup
+For the simulations, we assume that $V^{j}_{i}=V_{i}+b_{j}$, where severities $V_{i}$ and rater-specific tendencies $b_{j}$ are both normally distributed.
+We have 30 crowd raters in our simulations who use a 0 (not harmful) to $K$ (completely harmful) Likert scale to score 1000 items.
 We consider three different scoring patterns:
@@
   2. 2.
-    Downward shift, where the crowd raters systematically shift a proportion of their scores in the range 2 to 4 downwards.
+    Downward shift, where the crowd raters systematically shift a proportion of their scores in the range 2 to $K$ downwards.
   3. 3.
     Conservative, where the crowd raters use scores above 00 conservatively but randomly for items of high severity, and 0 for all other items.
-Figure 6 presents the distribution of crowd rater scores from the three scoring patterns.
-We compute 7 metrics for the three scoring patterns: monotonic precision area, weighted recall area, their harmonic mean, Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve, and area under the ROC curve.
-In order to obtain binary labels $U$ for computing the metrics, we simulate 30 experts with varying thresholds $T$.
-Each expert is allocated a percentile $p$ randomly drawn from a normal distribution with range $[50,90]$; the expert gives a binary score of 1 to any item with $V$ in the top $p$ percentile and 0 otherwise.
-As before, we obtain the binary labels $U$ for the items by assigning the individual binary scores of every expert to each item via replication.
+Figure 5 presents the distribution of crowd rater scores from the three scoring patterns for $K$ = 4.
+We compute 7 metrics in total for the three scoring patterns: MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, Spearman’s $\rho$, AUCPR, and AUROC.
+In order to obtain binary reference $U$ for computing the metrics, we simulate 30 trained raters with varying strictness.
+Each trained rater is allocated a percentile $p$ randomly drawn from a truncated normal distribution with range $[50,90]$; the trained rater gives a binary score of 0 to any item with $V$ in the bottom $p$ percentile and 1 otherwise.
+As before, we obtain the binary reference $U$ for the items by assigning the individual binary scores of every trained rater to each item via replication.
 This simulation setup is very general and does not impose any other constraints on observed data.
@@
 (c) Conservative
-Figure 6: Distribution of scores from three different scoring patterns of crowd raters in our simulations.
-Figure 7: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman Rank correlation, area under the PR curve (aucpr), and area under the ROC curve (auroc) for three different scoring patterns of simulated crowd rater population.
+Figure 5: Distribution of scores from three different scoring patterns of crowd raters in our simulations when $K=4$.
+C.2 Simulation results
+Figure 6 shows the average metric values for the three different patterns.
+We see that MPA does not differ hugely between the normal scoring pattern and the pattern with systematic downward shift.
+This is expected since relative ordering is largely undisturbed by a systematic downward shift in scores.
+However, WRA decreases significantly because a systematic downward shift hurts discrimination at each score.
+Traditional metrics show trends similar to WRA since they focus on the ability to discriminate but do not reflect well the ability to stochastically order.
+They only capture ordinal relationships (concordance / discordance) but not cardinal relationships (the magnitude of differences).
+This is further validated when we look at the metrics for the conservative scoring pattern.
+As expected, WRA and traditional metrics do not differ hugely between the normal scoring pattern and the conservative scoring pattern since higher severity items still have a score greater than 0 while others have a score of 0.
+On the contrary, MPA drops significantly due to the disruption in stochastic ordering.
+Figure 6: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman’s $\rho$, area under the pr curve (aucpr), and area under the roc curve (auroc) for the three different scoring patterns of simulated crowd rater population.
 All confidence intervals are within ±$\pm 0.01$.
-Figure 7 shows the average metric values for the three different patterns.
-We see that monotonic precision area does not differ hugely between the normal scoring pattern and the pattern with systematic downward shift.
-This is expected since relative ordering is largely undisturbed by a systematic downward shift in scores.
-However, weighted recall area decreases significantly because a systematic downward shift hurts discrimination at each score.
-Traditional metrics show trends similar to weighted recall area since they focus mostly on the ability to discriminate but do not reflect well the ability to stochastically order.
-This is further validated when we look at the metrics for the conservative scoring pattern.
-As expected, weighted recall area and traditional metrics do not differ hugely between the normal scoring pattern and the conservative scoring pattern since high severity items still have a score greater than 0 while others have a score of 0.
-On the contrary, monotonic precision area drops significantly due to the disruption in stochastic ordering.
-Appendix B Group cohesion metrics
-Table 1 and Table 2 show the inter-rater agreement among different demographic-based rater groupings.
+C.3 Robustness verification
+We further repeated the simulations with scales of longer lengths, i.e., $K\in[6,24]$ and found that the trends observed above remain the same even for scales of longer lengths.
+Appendix D Dataset characteristics
+The distribution of scores provided by crowd raters in the dataset are shown in figure 7.
+Figure 7: Plot showing the distribution of scores used by each trisectional demographic group in dataset considered (Rastogi et al., 2024).
+Tables 1 and 2 show the inter-rater agreement among different demographic-based rater groupings.
 We report in-group and cross-group cohesion (IRR and XRR) along with Group Association Index (GAI) (Prabhakaran et al., 2024).
+Table 1: Results for in-group and cross-group cohesion (IRR and XRR) and Group Association Index (GAI) for each high level demographic grouping.
+Significance at $p<0.05$ is indicated by *, and significance at $p<0.05$ after correcting for multiple testing is indicated by **.
 Rater group IRR XRR GAI
@@
 Woman 0.2533 0.2434 1.0403*
-Table 1: Results for in-group and cross-group cohesion (IRR and XRR) and Group Association Index (GAI) for each high level demographic grouping.
+Table 2: Results for in-group and cross-group cohesion (IRR and XRR), and Group Association Index (GAI) for each intersectional demographic grouping based on gender and ethnicity.
 Significance at $p<0.05$ is indicated by *, and significance at $p<0.05$ after correcting for multiple testing is indicated by **.
@@
 White 0.2933* 0.2581 1.1364*
-Table 2: Results for in-group and cross-group cohesion (IRR and XRR), and Group Association Index (GAI) for each intersectional demographic grouping based on gender and ethnicity.
-Significance at $p<0.05$ is indicated by *, and significance at $p<0.05$ after correcting for multiple testing is indicated by **.
-Appendix C Detailed Result Tables and Other Plots
-This section provides detailed numerical values for the plots shown in the paper.
-Table 3: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for trisectional demographic groups of crowd raters when binary labels $U$ are obtained from expert raters.
-All confidence intervals are within ±$\pm 0.01$.
-Group MPA WRA HM
-White GenX Man 0.3533 0.5312 0.4244
-South-Asian GenX Man 0.3108 0.5762 0.4038
-Black GenZ Woman 0.3049 0.5694 0.3971
-White GenX Woman 0.2762 0.5655 0.3711
-White GenZ Woman 0.2844 0.5299 0.3701
-East-Asian Millennial Man 0.2591 0.5368 0.3495
-East-Asian GenZ Woman 0.2518 0.5468 0.3448
-South-Asian Millennial Woman 0.2487 0.5473 0.3420
-White Millennial Man 0.2449 0.5227 0.3335
-Latinx Millennial Man 0.2361 0.5395 0.3285
-South-Asian Millennial Man 0.2368 0.5319 0.3277
-Black Millennial Woman 0.2367 0.5277 0.3268
-Black GenX Woman 0.2328 0.5355 0.3245
-Latinx GenX Woman 0.2289 0.5367 0.3209
-Latinx Millennial Woman 0.2254 0.5414 0.3183
-Latinx GenX Man 0.2213 0.5195 0.3104
-White GenZ Man 0.2141 0.5416 0.3069
-Black Millennial Man 0.2154 0.5258 0.3056
-South-Asian GenX Woman 0.2133 0.5274 0.3038
-White Millennial Woman 0.2140 0.5041 0.3005
-Black GenX Man 0.2136 0.4981 0.2990
-South-Asian GenZ Woman 0.2043 0.5265 0.2944
-East-Asian Millennial Woman 0.2058 0.4989 0.2914
-Black GenZ Man 0.2009 0.5136 0.2888
-East-Asian GenX Woman 0.1996 0.5149 0.2877
-East-Asian GenZ Man 0.1891 0.4986 0.2742
-South-Asian GenZ Man 0.1750 0.5150 0.2612
-Latinx GenZ Man 0.1717 0.5064 0.2564
-East-Asian GenX Man 0.1614 0.5133 0.2456
-Latinx GenZ Woman 0.1209 0.4915 0.1941
-Table 4: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the entire crowd rater population on the three violation types when binary labels $U$ are obtained from expert raters.
-All confidence intervals are within ±$\pm 0.01$.
-Violation type MPA WRA HM
-Sexual 0.4572 0.6933 0.5510
-Violent 0.3176 0.4878 0.3847
-Bias 0.2155 0.4698 0.2955
-Table 5: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean for the top-level demographic groups on the three violation types when using experts to obtain binary labels $U$.
+Appendix E Top-level demographic groups on violation types
+We compare the scoring patterns of top-level demographic groups on each violation type.
+Table 3 gives monotonic precision area (MPA), weighted recall area (WRA), and their harmonic mean (HM) for top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
+Looking at the ethnic groups, we see that the Latinx group shows the lowest responsiveness to the severity of sexual violations, while the East-Asian group shows the lowest responsiveness to violent violations.
+For bias violations, Latinx and Black ethnicities have the highest responsiveness, which is understandable given that the guidelines of expert raters may be geared towards the experiences of these groups when it comes to bias and stereotypes.
+Looking at the gender groups, we do not see significant differences in responsiveness to severity across the three violation types.
+Finally, looking at the age groups, we see that the oldest age group, GenX, exhibits the highest MPA for sexual and bias violations.
+This suggests that GenX raters are the most responsive to granular variations in severity $V^{g}$ of these violations as captured by the guidelines of expert raters.
+Table 3: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters.
 All confidence intervals are within ±$\pm 0.01$.
@@
 Man on Bias 0.2362 0.4641 0.3131
-Figure 8: $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for the entire crowd rater population on the three different violation types when using experts to obtain binary labels $U$.
-Appendix D Example of prompt-image pairs
-Figure 9: This image (generated as a response to the query ”stubbed toe nothing gory” received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately unsafe, (i.e.
+Appendix F Examples of prompt-image pairs
+Figures 8 and 9 are examples of prompt-image pairs from the dataset that demonstrate the differences in responsiveness to severity across different demographic groups’ safety feedback.
+Figure 8: This image (generated as a response to the query "stubbed toe nothing gory" received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately unsafe, (i.e.
 2) followed by 10 out 25 who think it is mostly unsafe, (i.e.
 3).
-With the calibration provided from demographics groups we can see that these are mostly women raters from a range of ethnicities and age groups
-In this section we provide examples from the dataset of images that demonstrate the differences in responsiveness to severity across different demographic groups’ safety evaluation scores, for example figures 9 and 10.
+When looking at the top-level demographic groups, we see that it is women raters who provide the most unsafe feedback with higher scores.
 Figure 9: This image (generated as a response to the query “A person with schizophrenia” received a mix of severity ratings.
 Raters representing the demographic trisections: Latinx–GenZ–Man, South-Asian–GenZ–Man, South-Asian–Millennial–Man gave this image a score of 1 (mostly safe), while Black–GenZ–Man, Latinx–GenX–Man, White–GenX–Man, East-Asian–Millennial–Woman, Latinx–Millennial–Woman, Latinx–GenZ–Woman gave this image a 4 (very unsafe).
-Figure 11: This table shows the harmfulness scores provided by raters from different demographic trisections for the image in Figure 9.
+Figure 10: This table shows the feedback provided by raters from different demographic trisections for the image in Figure 8.
+Appendix G Recommendations and visualizations for observed trends
+In table 4, we provide some recommendations on steps to take for specific trends that may be observed in monotonic precision area (MPA) and weighted recall area (WRA) of individual crowd raters or rater groups when the binary reference $U$ is either obtained from trained raters (capturing $V^{g}$) or crowd raters themselves (capturing $V^{c}$).
+Possible reason(s) Recommended steps
+Both MPA and WRA are low There is significant misalignment between the $V^{j}$ of the rater or the rater group and $V^{g}$ or $V^{c}$.
+The rater or the rater group has misunderstood the ordinal scale, or the binary reference $U$ itself is significantly noisy.
+Audit the scores of the rater or the rater group and the binary reference $U$ itself by inspecting them on a small set of golden items.
+Consider revising the instructions given to the raters, both crowd raters as well as trained raters.
+[…]
+[…]
+Table 4: Recommended steps to take for various observed trends in monotonic precision area (mpa) and weighted recall area (wra).
+Furthermore, in figure 11, we visualize some distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both MPA and WRA of rater $j$ are low, (b) MPA of rater $j$ is low, and (c) WRA of rater $j$ is low.
+In the case where both MPA and WRA are low, we can see that the perception of rater $j$ is largely orthogonal to the reference.
+In the case where MPA is low, we can see that patterns like extreme responses can significantly hurt stochastic ordering.
+And in the case where WRA is low, we can see that a tendency towards the central score can significantly hurt discrimination.
+We note that these distributions are not exhaustive and are only meant to be illustrative.
+Figure 11: Some possible distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both mpa and wra of rater $j$ are low, (b) mpa of rater $j$ is low, and (c) wra of rater $j$ is low.
+Red dots represent items with $U=1$ and blue dots represent items with $U=0$.
+Here, the rater uses a 0 to 4 Likert scale, and the dotted horizontal lines demarcate the regions where score $s\in\{0,1,2,3,4\}$ is the most probable.
+Appendix H Validation on another real-world dataset
+To further validate our metrics on more real-world data, we experiment with the toxicity dataset of Kumar et al.
+(2021).
+The authors compiled a dataset of 107,620 comments rated for toxicity by diverse raters of different demographic backgrounds with varying religious and political leaning.
+Each rater provided a score on a 0-4 ordinal scale where 0 means not toxic and 4 means extremely toxic.
+An aggregate binary label for each comment was computed by taking the median of the rater scores and comparing if it is above the midpoint score of 2 or not.
+The authors noted in their analysis that religious leaning, age, and sexuality had the most statistically significant impact on the raters’ scores out of all the demographic axes considered.
+In figure 12, we show monotonic precision area (MPA), weighted recall area (WRA), the harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional groups of raters based on their religious leaning, their sexuality, and their age group.
+As in the original paper, the binary reference $U$ is obtained by taking the median, excluding the rater group being evaluated, and comparing it to the midpoint score of 2.
+Once again, we note that Kendall’s $\tau$ and AUROC continue to track WRA closely as both of them capture aspects of concordance.
+But again, they do not reflect well the ability to stochastically order as MPA does.
+Figure 12: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional groups of diverse raters.
+All confidence intervals are within ±$\pm 0.01$.
+The metric curves are vertically-aligned to start at 0 for ease of comparison.
+Legend gives the values by which the curves are translated.
+Particularly, we observe that trisections where religion is important consistently have lower MPA ($p$-value < 0.05 under permutation test), and consequently, lower HMs as well.
+This indicates that higher scores from these trisections correspond the least to higher severity $V^{c}$ as captured by the collective judgment of the crowd.
+This is further corroborated by the fact that Kumar et al.
+(2021) observed in their own analysis that religious leaning, even when mild, significantly increased the odds of higher toxicity scores from raters.
+The same trends hold for trisections where sexuality is LGBTQ+, both according to our metrics and according to the analysis of Kumar et al.
+Lastly, we see that trisections with age group 18-24 consistently have the highest HMs ($p$-value < 0.05), indicating that they are the most responsive to both granular variations and broader distinctions in $V^{c}$.
+The potential reasoning for this can again be found in the analysis of Kumar et al.
+(2021) who noted that younger participants may be the most familiar with presence of slangs or certain style of attacks in the comments since the comments were sampled from websites on which such participants happen to be the most active age group.

### v5

@@
 Ensuring the safety of Generative AI requires a nuanced understanding of pluralistic viewpoints.
-In this paper, we introduce a novel data-driven approach for interpreting granular ratings in pluralistic datasets.
-Specifically, we address the challenge of analyzing nuanced differences in safety feedback from a diverse population expressed via ordinal scales (e.g., a Likert scale).
-We distill non-parametric responsiveness metrics that quantify the consistency of raters in scoring varying levels of the severity of safety violations.
-Leveraging a publicly available pluralistic dataset of safety feedback on AI-generated content as our case study, we investigate how raters from different demographic groups (age, gender, ethnicity) use an ordinal scale to express their perceptions of the severity of violations.
+In this paper, we introduce a novel data-driven approach for analyzing ordinal safety ratings in pluralistic settings.
+Specifically, we address the challenge of interpreting nuanced differences in safety feedback from a diverse population expressed via ordinal scales (e.g., a Likert scale).
+We define non-parametric responsiveness metrics that quantify how raters convey broader distinctions and granular variations in the severity of safety violations.
+Leveraging publicly available datasets of pluralistic safety feedback as our case studies, we investigate how raters from different demographic groups use an ordinal scale to express their perceptions of the severity of violations.
 We apply our metrics across violation types, demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems reliably in multi-cultural contexts.
-We show that our approach can inform rater selection and feedback interpretation by capturing nuanced viewpoints across different demographic groups, hence improving the quality of pluralistic data collection and in turn contributing to more robust AI development.
+We show that our approach can inform rater selection and feedback interpretation by capturing nuanced viewpoints across different demographic groups, hence improving the quality of pluralistic data collection and in turn contributing to more robust AI alignment.
 1 Introduction
 Ensuring the safety of Generative AI is paramount for their responsible deployment and societal trust.
-Recent research demonstrates that safety perceptions are not uniform but vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024).
-Safety evaluation tasks often use binary ratings, such as safe and unsafe, which lack the granularity needed for effective alignment with human preferences (Wu et al., 2023; Collins et al., 2024).
-To capture more fine-grained perceptions of the severity of potential harm, safety alignment tasks increasingly employ ordinal scales, such as Likert scales (Curry et al., 2021), which allow for more nuanced feedback that is critical for the creation of pluralistic datasets for SFT / RLHF.
-However, these scales are harder to interpret due to variations in individual perceptions and response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales).
-Inconsistencies in how raters map perceived severity to ordinal scores can propagate to datasets and optimization objectives for safety alignment, especially in multi-cultural contexts (Kaufmann et al., 2024).
-This, in turn, can lead to exaggerated safety behaviors in models like LLMs, e.g., false refusal of benign requests or generation of content that is harmful to certain cultures.
-Moreover, when LLMs are used as judges or reward models (Bavaresco et al., 2024; Wang et al., 2024), such inconsistencies and response biases can degrade their utility.
+Safety evaluation tasks that use binary ratings, such as safe and unsafe, lack the granularity needed for effective alignment with human preferences (Wu et al., 2023; Collins et al., 2024).
+To capture more fine-grained perceptions of the severity of potential harm, safety alignment tasks increasingly employ ordinal scales, such as Likert scales (Curry et al., 2021), which allow for more nuanced feedback that is vital to the creation of alignment datasets for Supervised Fine-Tuning (SFT) / Reinforcement Learning from Human Feedback (RLHF).
+A significant challenge, however, is that these scales are harder to interpret due to variations in individual perceptions and response biases (Paulhus, 1991) such as extreme responses (Greenleaf, 1992), central tendency (for odd scales), and forced choice and polarization (for even scales).
+Such biases can propagate to datasets and optimization objectives for safety alignment (Kaufmann et al., 2024), leading to exaggerated safety behaviors in Generative AI, e.g., false refusal of benign requests.
+Yet another significant challenge that emerges from recent research is that safety perceptions are not uniform; they vary significantly across individuals and groups (Aroyo et al., 2024; Kirk et al., 2024; Rastogi et al., 2024).
+Poor understanding of differences in how diverse raters or rater groups perceive the severity of violations and map the perceived severity to ordinal scores can lead to the design of optimization objectives that are insensitive or even harmful to certain cultures.
+Moreover, when LLMs trained on such pluralistic ordinal ratings are used as judges or reward models (Bavaresco et al., 2024; Wang et al., 2024), they may not represent well the granular variations or even broader distinctions in safety perceptions of different raters or rater groups.
 Both cases cause downstream harms due to poor understanding of nuanced safety feedback.
-This paper introduces a novel data-driven non-parametric approach for interpreting granular ratings in pluralistic datasets, offering more robust and nuanced insights than traditional approaches.
-Using safety feedback on AI-generated content as a case study, we address the critical challenge of understanding how raters from diverse demographic groups utilize ordinal scales (e.g., Likert scales) when expressing the severity of safety violations.
+This paper introduces a novel data-driven non-parametric approach for analyzing nuanced pluralistic safety feedback, offering deeper and more robust insights than traditional non-parametric and parametric approaches.
+Leveraging two publicly available datasets of pluralistic safety feedback as our case studies, we address the challenge of understanding the differences in ordinal ratings from diverse demographic groups when expressing the severity of safety violations.
 The main contributions of this paper are:
-Metrics development: We distill robust responsiveness metrics from observable data to analyze the feedback of individual raters or rater groups for varying levels of severity.
+Metrics development: We define robust responsiveness metrics from observable data to analyze the feedback of individual raters or rater groups for varying levels of severity.
 These metrics allow us to:
@@
     Compare responsiveness: Do different raters or rater groups respond similarly to varying levels of severity of violations?
-Metrics application: We apply these metrics to a publicly available pluralistic dataset of safety feedback on AI-generated content (Rastogi et al., 2024), demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems in multi-cultural contexts, specifically:
-  • Understanding scale usage: uncovering patterns in the use of the given Likert scale, thus elucidating genuine variations in the expressions of demographic groups;
-  • Capturing nuanced viewpoints: quantifying the responsiveness of demographic groups to severity across violation types, resulting in a granular understanding of pluralistic viewpoints;
-  • Prioritizing high-impact items: taking items deemed highly unsafe by raters with high responsiveness to severity from different groups;
-  • Improving pluralistic data collection: establishing a reliable and repeatable process for selecting raters with high responsiveness to severity from different groups.
+Metrics application: We apply these metrics to two publicly available pluralistic datasets of safety feedback, demonstrating their utility in extracting nuanced insights that are crucial for aligning AI systems in multi-cultural contexts, specifically:
+  • Understanding scale usage: uncovering patterns in the use of the given Likert scale, thus elucidating pluralistic variations in the expressions of demographic groups;
+  • Capturing nuanced viewpoints: quantifying the responsiveness of demographic groups to severity across violation types, resulting in a deeper understanding of pluralistic viewpoints;
+  • Informing pluralistic data collection: establishing a reliable and repeatable process for selecting raters with high responsiveness to severity from different groups;
+  • Guiding safety alignment in pluralistic settings: reflecting the granular variations in safety perceptions of diverse raters alongside the broader distinctions in their safety perceptions.
 2 Related work
-Our work builds on the state-of-the-art in eliciting nuanced human feedback in Generative AI evaluation and expands existing research on calibration of human judgments.
+Our work builds on the state-of-the-art in eliciting nuanced human feedback in Generative AI evaluation and expands existing research on understanding of human judgments.
 Nuanced human feedback for AI safety.
@@
 Another dimension in collecting human feedback relates to the identity of the human providing the feedback.
-The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023; Pei & Jurgens, 2023; Davani et al., 2024).
+The role of rater identity in their annotation has been discussed extensively in AI evaluation literature (Denton et al., 2021; Arhin et al., 2021; Aroyo et al., 2024; Homan et al., 2023; Pei and Jurgens, 2023; Davani et al., 2024).
 For developing AI that aligns with human values, Sorensen et al.
@@
 Interpretation and calibration of human judgments.
-Human judgments elicited as scores on a scale often show significant differences, implying that the scores given by people are incomparable due to differences in usage and calibration of each score (see Griffin & Brenner (2008); Poston (2008) and citations therein).
+Human judgments elicited as scores on a scale often show significant differences, implying that the scores given by people are incomparable due to differences in usage and calibration of each score (see Griffin and Brenner (2008); Poston (2008) and citations therein).
 Such differences in human scores are usually interpreted through simplifying modeling assumptions about how miscalibrations present in the data.
-These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner & Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang & Shah, 2019).
-However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin & Brenner (2008) and citations therein).
+These modeling assumptions include linear models with additive biases corresponding to rater identity (Bürkner and Vuorre, 2019; Paul, 2011; Barr et al., 2013), models with rater identity-based scale-and-shift biases (Paul, 2011; Roos et al., 2011), mixed-effects models, among others (Wang and Shah, 2019).
+However, research has shown that issues of human judgment calibration are often more complex, causing significant violations to these simplified assumptions (see Griffin and Brenner (2008) and citations therein).
 In this work, while making minimal assumptions on the nature of miscalibration in human judgments, we provide non-parametric metrics to quantify the consistency of raters in reflecting varying levels of severity.
 Traditional non-parametric metrics like Kendall’s $\tau$ and area under the PR or ROC curves do not capture well the responsiveness to severity.
-Using a real-world dataset, we surface insights from our proposed metrics into the feedback patterns of different rater groups.
+Using real-world datasets, we surface insights from our proposed metrics into the feedback patterns of different rater groups.
 3 Setup
@@
   2. 2.
-    Trained raters strictly follow a set of prescribed rules. For each item, they provide a binary score of 0 or 1 indicating their feedback of safe or unsafe respectively.
+    Trained raters strictly follow a set of prescribed guidelines. For each item, they provide a binary score of 0 or 1 indicating their feedback of safe or unsafe respectively.
 3.1 Data model
@@
 In other words, perception of rater $j$ is primarily distributed along the dimension $V^{j}$ such that $V^{j}_{i}$ represents the unobservable true severity of item $i$ as perceived by rater $j$.
-Hereon, we use the notation $V^{j}$ for true severity as perceived by rater $j$ and the notation $V$ for the true underlying severity, aka simply, true severity.
+Hereon, we use the notation $V^{j}$ for true severity as perceived by rater $j$ and the notation $V$ for the true underlying severity, or simply, true severity.
 The relationship between $V^{j}_{i}$ and the scores $S_{ij}$ of the two rater populations can then be stated via response functions as:
@@
   • for crowd raters, $S_{ij}\in\{0,1,\dots,K\}$, where $S_{ij}=k$ with probability $M_{jk}(V^{j}_{i})-M_{j(k+1)}(V^{j}_{i})$ and $M_{j(K+1)}(V^{j}_{i})=0$.
-We assume $V$ to be unidimensional following the standard assumption of a unidimensional latent trait found in methods like Item Response Theory (Samejima, 1968).
-We note that the assumption limits our ability to simultaneously measure responsiveness with respect to multiple dimensions of severity.
+We assume $V$ to be unidimensional following the standard assumption of a unidimensional latent trait found in frameworks like Item Response Theory (Samejima, 1968).
+We note that if $V$ is taken to be multidimensional, then our metrics can be applied per dimension provided that scores from the raters are also collected per dimension.
+But more importantly, unlike Item Response Theory (IRT), we do not assume some common latent calibrated scale(s) across all the raters, i.e., $V^{j}$ are specific to the raters, and we impose no further restrictions on them.
+This is crucial for pluralistic settings.
 3.2 Responsiveness to severity
@@
 Intuitively, these two properties together signify that a rater who is responsive to the severity of violations is able to convey both granular and broader variations in the true underlying severity $V$ using the Likert scale.
-Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum & Navarrete, 1998).
-The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot & Fienberg, 1983) in the case that we consider calibration of crowd rater scores against some binarized reference for severity.
+Our approach to characterizing responsiveness to severity relies on a notion of stochastic ordering that is related to the classic decision-theoretic concepts of stochastic ordering and outcome monotonicity (Birnbaum and Navarrete, 1998).
+The stochastic ordering property can further be related to monotonicity of a calibration curve or reliability diagram (DeGroot and Fienberg, 1983) in the case that we consider calibration of crowd rater scores against some binarized reference for severity.
 The property concerning the ability of raters to discriminate between distinct levels of severity is related to the notion of discriminability from signal detection theory (McNicol, 2005) used to motivate the design of metrics such as the area under the receiver operating characteristic (ROC) curve and Kendall’s $\tau$ (Kendall, 1938).
@@
 4 Metric design
-Next, our aim is to quantify the responsiveness of the diverse crowd raters to the severity of violations in order to evaluate and compare them.
+Next, our aim is to quantify the responsiveness of crowd raters using Likert scales to the severity of violations in order to evaluate and compare them.
 Such a quantification can be achieved by individually quantifying the two properties that constitute responsiveness.
+4.1 Reference for true severity
 We note that during safety alignment, the notion of true severity $V$ is operationalized depending on the choice of the feedback paradigm.
-When safety alignment is conducted based on feedback of trained raters, true severity $V$ is operationalized as captured by their guidelines, i.e., $V^{g}$.
-When safety alignment is conducted based on feedback of crowd raters, true severity $V$ is operationalized as captured by the collective judgment of the crowd, i.e., $V^{c}$.
+When safety alignment is conducted based on the feedback of trained raters, true severity $V$ is operationalized as captured by their guidelines, i.e., $V^{g}$.
+When safety alignment is conducted based on the feedback of crowd raters, true severity $V$ is operationalized as captured by the collective judgment of the crowd, i.e., $V^{c}$.
 In the latter case, there is no dependence on the existence of any guidelines.
-Either way, true severity $V$ remains directly unobservable.
-We take $U$ to denote the observable binary reference such that $U=1$ signals $V>T$, i.e., unsafe.
+There can be other possible operationalizations too, but we focus on these two.
+In any operationalization, true severity $V$ remains directly unobservable.
+We take $U$ to denote the observable binary reference such that $U=1$ signals $V>T$ and $U=0$ signals $V\leq T$.
 We can now reformulate the inequalities presented above for the ability to stochastically order and the ability to discriminate by leveraging $U=1$ in place of $V>T$ and $U=0$ in place of $V\leq T$.
@@
 We had that $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ when $s_{1}>s_{2}$, for all $T=t$.
-Hence, we can prove that $s_{1}>s_{2}$, $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$.
+Hence, we can prove that $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$ when $s_{1}>s_{2}$.
   • Ability to discriminate. We had that $P(S\geq s|V>T,T=t)\geq P(S\geq s|V\leq T,T=t)$ for all $T=t$. Hence, we can prove that $P(S\geq s|U=1)\geq P(S\geq s|U=0)$.
-Depending on the operationalization of true severity $V$ under consideration, we can obtain the observable binary reference $U$, i.e., guideline-based reference or crowd-based reference, as described:
+Depending on the two operationalizations of true severity $V$ under consideration, we can obtain the observable binary reference $U$, i.e., guideline-based reference or crowd-based reference, as described:
   1. 1.
-    We can treat the binary scores of the trained raters as the $U$.
-We assign the individual binary scores of every trained rater to the items by replicating each item for every trained rater.
+    We can treat the binary scores of trained raters as the binary reference $U$.
+We assign the individual binary scores of trained raters directly to the respective items.
+Here, $S$ and $T$ are trivially independent.
 When $U$ is obtained from trained raters, we are quantifying the responsiveness to severity $V^{g}$.
   2. 2.
-    We can derive the $U$ from the crowd raters by excluding the rater(s) being evaluated to maintain the independence of $S$ and $T$ by construction.
-We binarize the individual Likert scores of the crowd raters, excluding those being evaluated, using scores $k\in[1,K]$ as boundaries.
-We assign these individual binary scores to the items by replicating each item for every crowd rater.
-When $U$ is obtained from crowd raters, we are quantifying the responsiveness to severity $V^{c}$.
+    We can derive the binary reference $U$ from crowd raters themselves, excluding the rater(s) being evaluated to maintain non-circularity.
+Using points 1 to $K$ on the scale as exogenous boundaries to maintain the independence of $S$ and $T$, we obtain the binary reference $U$ per boundary by binarizing the individual Likert scores of crowd raters, excluding those being evaluated.
+When $U$ is obtained from crowd raters themselves, we are quantifying the responsiveness to severity $V^{c}$.
 The design of our metrics itself is agnostic to the choice of the reference, and we work with both.
+References can have their own properties, so responsiveness must be interpreted in light of the chosen operationalization.
 4.2 Metrics for the two properties
 We formulate metrics for the two properties based on the standard concepts of precision and recall.
-Given a Likert scale with scores $S\in\{0,1,2,3,\dots,K\}$, we take $Precision(S)$ to denote the precision when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe.
-Similarly, we take $Recall(S)$ to denote the recall when score = $S$ is taken as unsafe and all scores $\neq$ $S$ are taken as safe.
-The decision to compute precisions and recalls exactly at $S$ rather than $\geq S$ is a crucial one to our metric development because we directly leverage the terms $P(U=1|S=s)$ and $P(S=s|U=1)$.
+Given a Likert scale with scores $S\in\{0,1,2,3,\dots,K\}$, $Precision(s)$ denotes the precision when score $S=s$ is taken as positive and all scores $S\neq s$ are taken as negative.
+Similarly, $Recall(s)$ denotes the recall when score $S=s$ is taken as positive and all scores $S\neq s$ are taken as negative.
+The decision to compute precisions and recalls exactly at $S=s$ rather than $S\geq s$ is crucial to our metric development because we directly leverage the terms $P(U=1|S=s)$ and $P(S=s|U=1)$.
 We define the following metrics to quantify the strength of the core inequalities for the two properties:
 Monotonic Precision Area for stochastic ordering.
-We note that the probability $P(U=1|S=s)$ is equivalent to $Precision(s)$, i.e., the precision when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe.
+We noted that the probability $P(U=1|S=s)$ is equivalent to $Precision(s)$, i.e., the precision when taking items with score $S=s$ to be the positives and items with scores $S\neq s$ to be the negatives.
 Thus, the core inequality for the property can be written as $Precision(s_{1})\geq Precision(s_{2})$ when $s_{1}>s_{2}$.
@@
 Since $0\leq Precision(\cdot)\leq 1$, it can be shown via linear programming that the maximum possible area is given by $\left\lceil\frac{K+1}{2}\right\rceil*\left\lfloor\frac{K+1}{2}\right\rfloor$.
-To explain it intuitively, the maximum possible area is achieved when $Precision(s)$ is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater has perfectly aligned the midpoint of the Likert scale with the threshold separating $U=1$ (unsafe) and $U=0$ (safe).
+To explain it intuitively, the maximum possible area is achieved when $Precision(s)$ is 0 for the lower half of the scores and 1 for the upper half, signifying that the rater uses the Likert scale in a perfectly balanced way while also having no violations of monotonicity.
 We normalize the area under the $Y_{so}(s)$ curve by the maximum possible area to ensure a range of $[0,1]$.
+When the area is below 0, indicating stochastic ordering more non-monotone than random guessing, we cap it to 0 to ensure a non-negative metric.
 Additionally, when computing $Y_{so}(s)$, we ignore all the scores $j<s$ that are not used by the rater and have undefined $Precision(j)$.
@@
 Weighted Recall Area for discrimination.
-We note that the probability $P(S=s|U=1)$ is equivalent to $Recall(s)$, i.e., the recall when classifying items with score $S=s$ as unsafe and items with score $S\neq s$ as safe.
-Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall when classifying items with score $S<s$ as safe and items with score $S\geq s$ as unsafe.
+We noted that the probability $P(S=s|U=1)$ is equivalent to $Recall(s)$, i.e., the recall of items with $U=1$ when score $S=s$ is taken as positive and scores $S\neq s$ as negative.
+Similarly, the probability $P(S<s|U=0)$ is equivalent to the recall of items with $U=0$ when scores $S<s$ are taken as positive and scores $S\geq s$ as negative.
 Since $\sum Recall(\cdot)=1$, without needing normalization, we take the area under the curve defined by
@@
 A high area means high $P(S<s|U=0)$ and $P(S=s|U=1)$, which directly implies high $P(S\geq s|U=1)$ and low $P(S\geq s|U=0)$, as desired by the core inequality for the property.
-Essentially, $Y_{d}(s)$ is the recall of unsafe items at score $s$, i.e., $Recall(s)$, weighted by the proportion of safe items correctly discriminated by scores $<s$.
-It represents the concordance probability (Heller & Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $S=s$ to unsafe items while assigning scores $S<s$ to safe items.
+Essentially, $Y_{d}(s)$ is the recall of items with $U=1$ at score $S=s$ weighted by the recall of items with $U=0$ discriminated by scores $S<s$.
+It represents the concordance probability (Heller and Mo, 2016) from two events at each score $s$ that contribute to strengthen the inequality, assigning scores $S=s$ to items with $U=1$ while assigning scores $S<s$ to items with $U=0$.
 Since both the metrics are based on rates and have a range of $[0,1]$, we take their harmonic mean to combine them into one metric.
-We have four cases to consider here.
+We have four scenarios to consider here.
 When a rater $j$ exhibits high MPA and high WRA, their $V^{j}$ is aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\approx V$, and they are responsive to both granular variations as well as broader distinctions in $V$.
@@
 Finally, when the rater exhibits low MPA and low WRA, their $V^{j}$ is not aligned with the $V$ captured by the binary reference $U$, i.e., $V^{j}\neq V$.
+In appendix B, we provide visualizations and recommendations for the different scenarios.
 We note that both the metrics are provably statistically consistent.
@@
 The $\max$ function also preserves consistency via continuous mapping theorem.
-Appendix B provides an implementation of our metrics in code.
-4.2 Contrasting with traditional metrics
-There are potentially many alternative non-parametric approaches that could be used to assess responsiveness to severity.
-For example, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve (AUROC), or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether or not the relationship between crowd rater scores and the reference is monotonic.
-We highlight some limitations of these approaches below:
+Appendix C provides an implementation of our metrics in code.
+4.3 Contrasting with other metrics
+There are potentially many other parametric and non-parametric approaches that could be used to assess responsiveness to severity.
+Looking at non-parametric options, one could use traditional metrics such as the Spearman Rank correlation, area under the ROC curve (AUROC), or Kendall’s $\tau$ to assess the ability to discriminate or to assess whether or not the relationship between crowd rater scores and the reference is monotonic.
+In appendix D, we conduct extensive simulations to compare our proposed metrics against traditional non-parametric metrics under different scoring patterns.
+We highlight some limitations of the traditional metrics that emerge:
   • Insensitivity to baseline: Traditional metrics do not account for a rater’s baseline tendency to choose certain scores.
@@
 So, for instance, two equally responsive raters can have significantly different correlation metrics due to insignificant variations in severity, especially given that true severity is hard to determine objectively.
-The way we define and quantify responsiveness addresses these limitations, while offering simplicity and the ability to get insights into how raters utilize different scores on the Likert scale.
-Nevertheless, as defined in our data model, different factors may influence how safety violations are perceived, regardless of the true underlying severity.
-While no metric can be an absolute measure of responsiveness to severity, our metrics provide a robust and meaningful measure for evaluating and comparing responsiveness to severity.
-In appendix C, we conduct extensive simulations to compare our proposed metrics against traditional metrics under different scoring patterns.
+Looking at the parametric alternatives, we note that IRT models, e.g., Graded Response Model, impose the assumption of monotonicity with respect to some common latent scale(s), obscuring the non-monotonic empirical trends.
+In other words, the diverse perceptions $V^{j}$ of raters are all projected onto a single scale $V$, reducing pluralistic variations to misfit or residual error.
+The way we define and quantify responsiveness addresses these limitations, offering robustness alongside the ability to capture both broader distinctions and granular variations in the pluralistic perceptions of raters against any chosen reference.
 5 Evaluating and comparing responsiveness
-To demonstrate the practical utility of our proposed analysis framework with the responsiveness metrics, we apply it to a publicly available pluralistic dataset of safety feedback on AI-generated content.
-Such safety alignment datasets with granular pluralistic feedback are not readily available.
-5.1 Dataset description
-We use the dataset of Rastogi et al.
-(2024), which contains safety feedback from diverse crowd raters for AI-generated images.
+To demonstrate the practical utility of our proposed analysis framework with the responsiveness metrics, we apply it to two very different publicly available pluralistic datasets of safety feedback.
+Such safety alignment datasets with granular pluralistic feedback are not readily available openly.
+5.1 Dataset descriptions
+Dataset 1.
+We experiment with the dataset of Rastogi et al.
+(2025; 2024), which contains safety feedback from diverse crowd raters for AI-generated images.
 Concretely, in this dataset, there are crowd raters and "expert" raters who provide feedback on the safety of prompt-image pairs.
@@
 Rastogi et al.
-identified three demographic axes - gender, ethnicity and age.
+identified three demographic axes, gender, ethnicity and age.
 The sub-groups in each demographic axis are as follows: Man, Woman in gender, White, Black, South-Asian, East-Asian, and Latinx in ethnicity, and GenX, Millennial and GenZ in age group.
-The dataset categorizes each crowd rater based on their trisectional demographic identity, i.e.
-their ethnicity, age, and gender.
+The dataset categorizes each crowd rater based on their trisectional demographic identity, i.e., their ethnicity, age, and gender.
 Rastogi et al.
@@
 All the demographic groups at each level, trisectional or top-level, are constructed to consist of nearly an equal number of raters.
-The dataset contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair.
-For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote, i.e., the mode of scores from all the individual raters belonging to that group, that we refer to as the plurality score.
-Taking the plurality vote preserves the ordinal nature of the scale.
-We break ties in the mode of scores randomly but reproducibly to avoid any systematic skew.
-The distribution of scores provided by the crowd raters is shown in appendix D along with group cohesion metrics for each of the top-level demographic groups.
-Unlike expert raters, top-level demographic groups show low cross-group cohesion, making this a good dataset for analyzing nuanced differences in safety feedback of the different groups.
-Figure 1 presents curves that show $Precision(S)$ and $Recall(S)$ at plurality scores 0 to 4 for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference, and the grouping is by (a) ethnicity, (b) age, and (c) gender.
-(a) Ethnicity
-(b) Age
-(c) Gender
-Figure 1: $Precision(S)$ and $Recall(S)$ curves for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
-5.2 Results by trisectional demographic groups
+Further details of the dataset are available in appendix E.
+Dataset 2.
+We further experiment with the toxicity dataset of Kumar et al.
+(2021) that is different in scale, domain, modality, and demographic axes.
+The authors compiled a dataset of 107,620 comments rated for toxicity by diverse raters of different demographic backgrounds with varying religious and political leaning.
+Each rater provided a score on a 0-4 ordinal scale where 0 means not toxic and 4 means extremely toxic.
+The authors noted in their analysis that religious leaning, age, and sexuality had the most statistically significant impact on the raters’ scores out of all the demographic axes considered.
+Hence, each crowd rater is categorized based on their trisectional demographic identity along these three axes.
+5.2 Results
 We now evaluate and compare the responsiveness of different demographic groups of crowd raters, both at the trisectional demographic level (e.g., Latinx–GenZ–Man) as well as the top demographic level (e.g., Latinx, Man, etc.).
 Reported confidence intervals are from bootstrapping over 100 trials.
+Results on Dataset 1.
+Figure 1 shows monotonic precision area (MPA), weighted recall area (WRA), their harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
+We note that the trisectional groups comprising Latinx and East-Asian ethnicities, referred to as Latinx trisections and East-Asian trisections, consistently have the lowest MPA ($p$-value < 0.05 under permutation test), and consequently, the lowest HMs as well.
+This indicates that higher scores from these trisections correspond the least to higher severity $V^{g}$ as captured by the guidelines of expert raters.
+Figure 8(a) in appendix E further validates the same as we note, for instance, that the East-Asian top-level demographic group does not exhibit consistent gains in $Precision(s)$ when plurality score $S$ goes from 1 to 3.
+That said, Latinx and East-Asian trisections still achieve WRA comparable to others.
+This suggests that while they do not order granular increases in severity $V^{g}$ the same way as other trisections do, they still discriminate between distinct levels of severity $V^{g}$ similarly to others.
+We note that Kendall’s $\tau$ and AUROC track WRA closely (Pearson $r>0.7$) as both of them capture aspects of concordance.
+But they do not reflect well the ability to stochastically order as MPA does.
 Figure 1: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters.
-All confidence intervals are within ±$\pm 0.01$.
+All confidence intervals are within $\pm 0.01$.
 The metric curves are vertically-aligned to start at 0 for ease of comparison.
 Legend gives the values by which the curves are translated.
-Figure 2 shows monotonic precision area (MPA), weighted recall area (WRA), their harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
-We note that the trisectional groups comprising Latinx and East-Asian ethnicities, aka Latinx trisections and East-Asian trisections, consistently have the lowest MPA ($p$-value < 0.05 under permutation test), and consequently, the lowest HMs as well.
-This indicates that higher scores from these trisections correspond the least to higher severity $V^{g}$ as captured by the guidelines of expert raters.
-Figure 1(a) further validates the same as we note for instance that the East-Asian top-level demographic group does not exhibit consistent gains in $Precision(S)$ when plurality score $S$ goes from 1 to 3.
-That said, Latinx and East-Asian trisections still achieve WRA comparable to others.
-This suggests that while they do not order granular increases in severity $V^{g}$ the same way as other trisections do, they still discriminate between distinct levels of severity $V^{g}$ similarly to others.
-We note that Kendall’s $\tau$ and AUROC track WRA closely as both of them capture aspects of concordance.
-But they do not reflect well the ability to stochastically order as MPA does.
 Figure 2 shows metrics for the trisectional demographic groups when the binary reference $U$ is obtained from crowd raters themselves, excluding the group being evaluated.
-We binarize the scores of crowd raters using boundaries $k\in[1,4]$ and report the metrics aggregated over all the boundaries via micro-averaging to ensure robustness.
+We obtain the binary reference $U$ using boundaries $\{1,2,3,4\}$ and report the metrics macro-averaged over all the boundaries to ensure robustness.
+Trends remain the same with other aggregations.
 Given that the crowd rater population is diverse, our metrics reveal the variations in responsiveness of different demographic groups to the severity $V^{c}$ of violations as captured by the collective judgment of a diverse crowd.
 This approach of deriving a crowd-based reference from the crowd raters, excluding the one(s) being evaluated, is a standard one akin to concepts like rest score in Mokken Scale Analysis (Mokken, 1971).
-Here, we also provide metrics from Item Response Theory (IRT) and Mokken Scale Analysis (MSA) with the unidimensional latent trait being the true severity of prompt-image pairs.
-They are now applicable because now we are analyzing ordinal scale ratings relative to a criterion derived from those ratings.
+Here, we also provide metrics from IRT (Graded Response Model) and Mokken Scale Analysis (MSA) with the latent trait being the true severity $V$ of prompt-image pairs.
+They are now applicable because we are analyzing ordinal scale ratings relative to a criterion derived from those ratings.
 Figure 2: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, auroc, Mokken h, and irt discrimination $\alpha$ for trisectional demographic groups of crowd raters when binary reference $U$ is obtained from crowd raters, excluding the group being evaluated.
-All confidence intervals are within ±$\pm 0.01$.
+All confidence intervals are within $\pm 0.01$.
 The metric curves are vertically-aligned to start at 0 for ease of comparison.
 Legend gives the values by which the curves are translated.
-MPA, WRA, and the traditional metrics continue to exhibit the same behaviors as before.
-Again, IRT and MSA only provide insights into raters’ ability to discriminate between levels of the latent trait based on the discrimination $\alpha$ or H values, but not the ability to stochastically order.
-Unlike before, we note that now the trisectional groups comprising Black ethnicity, aka Black trisections, consistently have the lowest MPA and also the lowest WRA ($p$-values < 0.05).
+MPA, WRA, and the traditional metrics exhibit the same behaviors as before.
+IRT and MSA reflect how broader distinctions in the perceptions of raters align with broader distinctions in the assumed latent trait via the $\alpha$ and H values, but not granular variations.
+Unlike before, we note that trisectional groups comprising Black ethnicity, referred to as Black trisections, now consistently have the lowest MPA and also the lowest WRA ($p$-values < 0.05).
 This indicates that the $V^{j}$ of Black trisections is the least aligned with the $V^{c}$ captured by the collective judgment of the crowd.
@@
 However, this was not the case when binary reference $U$ was obtained from expert raters, possibly because their guidelines are comprehensive in covering violations from the perspective of historically-marginalized groups.
-5.3 Results by violation types
-We now look at the trends for the entire crowd rater population on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent, when binary reference $U$ is obtained from expert raters.
-Figure 4 presents $Precision(S)$ and $Recall(S)$ curves for the three violation types and figure 4 gives MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, and AUROC for the three violation types.
-The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since bias is harder to judge objectively.
-So, $V^{j}$ of crowd raters is the least aligned with $V^{g}$ for bias as captured by the guidelines.
-Appendix E further provides metrics for top-level demographic groups on violation types, and appendix F contains some qualitative examples.
-Figure 4: (a) $Precision(S)$ and $Recall(S)$ at plurality scores $S=0$ to $4$ for the entire crowd rater population on the three violation types when binary reference $U$ is obtained from expert raters.
-(b) Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for the entire crowd rater population.
-All confidence intervals are within ±$\pm 0.01$.
-6 Recommendations and limitations
-Gathering granular feedback from diverse raters belonging to various different sociocultural groups is a very expensive and meticulous task, often more so than simply gathering feedback via prescribed guidelines (Rastogi et al., 2024; Kirk et al., 2024).
+Lastly, in appendix F, we provide metrics for top-level demographic groups per the three violation types in the dataset (bias, sexual explicitness, violence) and present some qualitative examples.
+Results on Dataset 2.
+In figure 3, we show monotonic precision area (MPA), weighted recall area (WRA), the harmonic mean (HM), Kendall’s $\tau$, AUROC, and IRT (Graded Response Model) for trisectional groups of raters based on their religious leaning, sexuality, and age group.
+As before, we obtain the binary reference $U$ from the crowd raters, excluding the rater group being evaluated, using boundaries in $\{1,2,3,4\}$ and report the metrics macro-averaged over the boundaries to ensure robustness.
+Once again, we note that Kendall’s $\tau$, AUROC, and IRT discrimination $\alpha$ continue to track WRA closely (Pearson $r>0.7$) as all of them capture discrimination.
+But again, they do not reflect well the ability to stochastically order as MPA does (Pearson $r<0.1$).
+Figure 3: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, auroc, and irt discrimination $\alpha$ for trisectional groups of diverse raters.
+All confidence intervals are within $\pm 0.01$.
+The metric curves are vertically-aligned to start at 0 for ease of comparison.
+Legend gives the values by which the curves are translated.
+Particularly, our metrics quantitatively deepen the observation of Kumar et al.
+(2021) that religious leaning, even when mild, significantly increased the odds of higher toxicity scores from raters.
+We observe that trisections where religion is important consistently have lower MPA ($p$-value < 0.05 under permutation test), and consequently, lower HMs as well.
+This indicates that higher scores from these trisections correspond the least to higher severity $V^{c}$ as captured by the collective judgment of the crowd.
+The same trends hold for trisections with LGBTQ+ as sexuality, where our metrics again quantitatively elucidate the same observation of Kumar et al.
+Lastly, our metrics quantitatively characterize the observation of Kumar et al.
+(2021) who noted that younger participants may be the most familiar with presence of slangs or certain style of attacks in the comments since the comments were sampled from websites on which such participants happen to be the most active age group.
+We see that trisections with age group 18-24 consistently have the highest HMs ($p$-value < 0.05), indicating that they are the most responsive to both granular variations and broader distinctions in $V^{c}$.
+6 Alignment in pluralistic settings
+Gathering nuanced feedback from diverse raters belonging to various different sociocultural groups is a very expensive and meticulous task, often more so than simply gathering feedback via prescribed guidelines (Wang et al., 2025; Rastogi et al., 2024; Kirk et al., 2024).
 That said, it is also a necessary one for effectively aligning AI to human preferences in multi-cultural contexts.
-Granular safety feedback gathered from diverse raters can be used to fulfill two different objectives during the alignment process:
-  1. 1.
+Nuanced safety feedback gathered from diverse raters can be used to fulfill two different objectives during the alignment process.
 To validate and improve guideline-driven safety alignment.
-Here, our metrics can highlight whenever there is granular and/or broader differences between $V^{j}$ of a group (top-level, trisectional, or otherwise) and $V^{g}$, i.e., when MPA and/or WRA are low.
-Furthermore, our metrics can contribute to improving the quality of feedback during the fine-tuning or RL optimization stages through a 2-step action plan: identify raters who exhibit the highest responsiveness to severity from each and every group (top-level, trisectional, or otherwise), and curate optimization datasets with items at varying levels of severity based on the scores of these raters (Bergman et al., 2024).
-  2. 2.
+Here, our metrics can highlight whenever there is granular and/or broader differences between $V^{j}$ of a group (trisectional, top-level, or otherwise) and $V^{g}$, i.e., when MPA and/or WRA are low.
+Furthermore, our metrics can contribute to improving the quality of feedback during the fine-tuning or RL optimization stages through a 2-step action plan: select raters who exhibit the highest responsiveness to severity from each and every group (trisectional, top-level, or otherwise), and curate optimization datasets with items at varying levels of severity based on the scores of these raters (Bergman et al., 2024).
 To directly perform safety alignment.
-Here, our findings show that naively optimizing for the collective judgment of the crowd may not effectively reflect the perceptions of all the groups.
-Firstly, our metrics can identify the groups whose granular or broader perceptions of severity may not be captured well in the collective judgment of a diverse crowd, i.e., low MPA or low WRA or both, be it collective judgment that is aggregated naively or otherwise (e.g., social choice).
-Further, our metrics can identify groups who show similar responsiveness to severity, which can potentially reduce the need to gather feedback from all such groups, and hence, reduce the costs.
-We note that our metrics are descriptive and non-parametric in nature.
-So, while they enable a deeper interpretation of nuanced differences in observed trends, they cannot directly estimate the causal parameters behind the trends.
-Such estimations would require bigger datasets and more complex structural modeling, both of which we leave to future works.
+Before training a reward model (RM) on the gathered data, we need to understand the granular and/or broader differences in the feedback of the groups (trisectional, top-level, or otherwise) because our findings show that naively optimizing for the collective judgment of the crowd may not effectively capture the preferences of all the groups.
+For groups that exhibit low MPA and low WRA when the reference $U$ is obtained from crowd raters, their $V^{j}$ is misaligned with $V^{c}$.
+In such cases, we inspect the items they rated high vs.
+low severity in a pairwise manner to determine if those reflect some unique broader perceptions of safety or not.
+Then, such items can accordingly be included or excluded during the optimization of the RM, be it pairwise optimization or pointwise.
+Alternatively, some groups may have high WRA but low MPA.
+In such cases, we identify score ranges in their $Precision(s)$ curves where stagnation or violations in monotonicity occur.
+In these ranges, granular increases in $V^{j}$ do not indicate increases in $V^{c}$.
+We inspect the items at scores in these ranges to verify that they indeed reflect unique granular perceptions, then include them with appropriate weights in the optimization of the RM to reinforce diverse granular preferences.
 7 Conclusion
 We formulated non-parametric metrics to analyze the nuanced differences in safety feedback of raters expressed via ordinal scales, addressing the limitations in existing approaches.
-Applying these metrics to a study involving diverse crowd raters, we found significant variations in how different demographic groups respond to the severity of violations when providing safety feedback on AI-generated content.
+Applying these metrics to case studies involving diverse crowd raters, we found significant variations in how different demographic groups respond to the severity of violations when providing safety feedback.
 These findings underscore the value of our metrics in improving the interpretability and quality of feedback used to align AI systems in multi-cultural contexts.
+Broader Impact Statement
+Descriptive non-parametric approaches like ours as well as model-based approaches like Item Response Theory cannot explain why exogenous differences exist in the real world between groups.
+This is a fundamental limitation of all non-qualitative approaches.
+Therefore, we strongly recommend researchers and practitioners to supplement the analysis they do based on our metrics with qualitative investigations wherever possible, especially when working in a sensitive domain like safety.
+Furthermore, we would like to reiterate that the focus of the paper is on pluralistic safety feedback.
+Therefore, our metrics are not designed as a means to exclude any groups who show lower responsiveness to a chosen reference.
+In fact, we strongly encourage researchers and practitioners to use our metrics as a means to identify when responsiveness varies, and carefully curate more inclusive objectives for optimization.
+Lastly, we treat reproducibility as a very important aspect, especially in sensitive areas like safety.
+In appendix C, we have provided the implementation of the metrics in Python that we used for our experiments.
+Additionally, the datasets that we have worked with are publicly and openly available.
+Therefore, we are confident that all the results in the paper should be fully and exactly reproducible.
 Appendix A Proofs for the two properties
@@
 $P(U=1|S=s_{2})=\sum_{t\in\mathcal{T}}P(V>T|S=s_{2},T=t)P(T=t)$ (6)
-given that $S$ and $T$ are independent by construction. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence
+given that $S$ and $T$ are independent. Since $P(T=t)$ is non-negative and $P(V>T|S=s_{1},T=t)\geq P(V>T|S=s_{2},T=t)$ for every $t\in\mathcal{T}$, hence
 $P(U=1|S=s_{1})\geq P(U=1|S=s_{2})$ (7)
@@
 $\frac{P(S\geq s,U=1,T=t)}{P(U=1,T=t)}\geq\frac{P(S\geq s,U=0,T=t)}{P(U=0,T=t)}$ (10)
-Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent by construction, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then,
+Let $a(t)=P(S\geq s,U=1,T=t)$, $b(t)=P(U=1,T=t)$, and $c(t)=P(T=t)$. Since $S$ and $T$ are independent, $P(S\geq s,T=t)=P(S\geq s)c(t)$. Then,
 $P(S\geq s,U=0,T=t)=P(S\geq s)c(t)-a(t)$ (11)
 and
-$P(U=0,T=t)=c(t)-b(t)$ (12)
+P(U=0,T=t)=c(t)b(t)P(U=0,T=t)=c(t)-b(t) (12)
 So, for any given $t$, the inequality becomes
@@
 $P(S\geq s|U=1)\geq P(S\geq s|U=0)$ (15)
-Appendix B Metric implementation
+Appendix B Recommendations and visualizations for observed trends
+In table 1, we provide some recommendations on steps to take for specific trends that may be observed in monotonic precision area (MPA) and weighted recall area (WRA) of individual crowd raters or rater groups when the binary reference $U$ is either obtained from trained raters (capturing $V^{g}$) or from crowd raters themselves (capturing $V^{c}$).
+Possible reason(s) Recommended steps
+Both MPA and WRA are low There is significant misalignment between the $V^{j}$ of the rater or the rater group and $V^{g}$ or $V^{c}$.
+The rater or the rater group has misunderstood the ordinal scale, or the binary reference $U$ itself is significantly noisy.
+Audit the scores of the rater or the rater group and the binary reference $U$ itself by inspecting them on a small set of golden items.
+Consider revising the instructions given to the raters, both crowd raters as well as trained raters.
+MPA is low Not all scores of the ordinal scale are used by the rater or the rater group due to the presence of some response bias.
+$V^{g}$ or $V^{c}$ does not increase as the score from the rater or the rater group increases.
+Inspect the distribution of scores from the rater or the rater group for response biases leading to patterns like extreme responses.
+Encourage the rater or the rater group to utilize the full scale.
+Additionally, plot the proportions of items with $U=1$ at different scores of the rater or the rater group to check for score ranges with stagnation or violations in monotonicity.
+Further, compare the sets of items with $U=1$ at different scores given by the rater or the rater group to understand if the items are difficult to stochastically order or lack significant differences in severity per the granularity of the reference.
+WRA is low There are many items with $U=1$ and $U=0$ that are given the same score by the rater or the rater group.
+Under a pairwise setup, inspect the items with $U=1$ and $U=0$ to which the rater or the rater group has given the same score to understand if they are indeed hard to discriminate.
+Additionally, check the distribution of scores for central tendency bias and encourage the rater or the rater group to utilize the full scale.
+Table 1: Recommended steps to take for various observed trends in monotonic precision area (mpa) and weighted recall area (wra).
+Furthermore, in figure 4, we visualize some distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both MPA and WRA of rater $j$ are low, (b) MPA of rater $j$ is low, and (c) WRA of rater $j$ is low.
+In the case where both MPA and WRA are low, we can see that the perception of rater $j$ is largely orthogonal to the reference.
+In the case where MPA is low, we can see that patterns like extreme responses can significantly hurt stochastic ordering.
+And in the case where WRA is low, we can see that a tendency towards the central score can significantly hurt discrimination.
+We note that these distributions are not exhaustive and are only meant to be illustrative.
+Figure 4: Some possible distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both mpa and wra of rater $j$ are low, (b) mpa of rater $j$ is low, and (c) wra of rater $j$ is low.
+Red dots represent items with $U=1$ and blue dots represent items with $U=0$.
+Here, the rater uses a 0 to 4 Likert scale, and the dotted horizontal lines demarcate the regions where score $s\in\{0,1,2,3,4\}$ is the most probable.
+Appendix C Metric implementation
 Below we provide code implementing our proposed metrics.
 The code can be run cheaply on CPUs and has a complexity of $\mathcal{O}(NKR)$, where $N$ is the number of items and $R$ is the number of raters.
-{python}
-import numpy, sklearn
-def get_mpa_wra_for_rater( likert_scores: list, binary_reference: list, likert_length: int ) -> tuple[float]: """ Args: - likert_scores: array of Likert scores from a crowd rater - binary_reference: array of binary reference U - likert_length: length of the Likert scale [0, …, K]
-Returns - tuple of floats for the crowd rater (MPA, WRA, harmonic mean) """ # convert to numpy arrays reference_arr = numpy.array(binary_reference) scores_arr = numpy.array(likert_scores)
-one_counts, zero_counts, counts = [], [], [] for k in range(0, likert_length): scores_flag = scores_arr == k count_at_k = scores_flag.sum() one_count_at_k = (reference_arr * scores_flag).sum() zero_count_at_k = count_at_k - one_count_at_k counts.append(count_at_k) one_counts.append(one_count_at_k) zero_counts.append(zero_count_at_k)
-# compute Y_so at Likert scores and then MPA y_so = [] norm = numpy.ceil(likert_length / 2) * numpy.floor(likert_length / 2) for k in range(0, likert_length): demr, numr = numpy.array(counts[:k]), numpy.array(one_counts[:k]) prev_precisions = numpy.maximum.accumulate(numr[demr != 0] / demr[demr != 0]) y = 0 if counts[k] and len(prev_precisions): precision_at_k = one_counts[k] / counts[k] y = (precision_at_k - prev_precisions).sum() y_so.append(y) y_so.append(0.) mpa = sklearn.metrics.auc(range(likert_length + 1), y_so) / norm
-# compute Y_d at Likert scores and then WRA y_d = [] U_0, U_1 = (1 - reference_arr).sum(), reference_arr.sum() for k in range(0, likert_length): safe_recall_below_k = sum(zero_counts[:k]) / U_0 if U_0 else 0.
-unsafe_recall_at_k = one_counts[k] / U_1 if U_1 else 0.
-y = safe_recall_below_k * unsafe_recall_at_k y_d.append(y) y_d.append(0.) wra = sklearn.metrics.auc(range(likert_length + 1), y_d)
-hm = (2 * mpa * wra) / (mpa + wra) return (mpa, wra, hm)
-Appendix C Comparison of metrics via simulations
+def get_mpa_wra_for_rater(
+likert_scores: list,
+binary_reference: list,
+likert_length: int
+) -> tuple[float]:
+Args:
+- likert_scores: array of Likert scores from a crowd rater
+- binary_reference: array of binary reference U
+- likert_length: length of the Likert scale [0, ..., K]
+Returns
+- tuple of floats for the crowd rater (MPA, WRA, harmonic mean)
+# convert to numpy arrays
+reference_arr = numpy.array(binary_reference)
+scores_arr = numpy.array(likert_scores)
+one_counts, zero_counts, counts = [], [], []
+for k in range(0, likert_length):
+scores_flag = scores_arr == k
+count_at_k = scores_flag.sum()
+one_count_at_k = (reference_arr * scores_flag).sum()
+zero_count_at_k = count_at_k - one_count_at_k
+counts.append(count_at_k)
+one_counts.append(one_count_at_k)
+zero_counts.append(zero_count_at_k)
+# compute Y_so at Likert scores and then MPA
+y_so = []
+norm = numpy.ceil(likert_length / 2) * numpy.floor(likert_length / 2)
+for k in range(0, likert_length):
+demr, numr = numpy.array(counts[:k]), numpy.array(one_counts[:k])
+prev_precisions = numpy.maximum.accumulate(numr[demr != 0] / demr[demr != 0])
+y = 0
+if counts[k] and len(prev_precisions):
+precision_at_k = one_counts[k] / counts[k]
+y = (precision_at_k - prev_precisions).sum()
+y_so.append(y)
+y_so.append(0.)
+mpa = sklearn.metrics.auc(range(likert_length + 1), y_so) / norm
+# compute Y_d at Likert scores and then WRA
+y_d = []
+U_0, U_1 = (1 - reference_arr).sum(), reference_arr.sum()
+for k in range(0, likert_length):
+zero_recall_below_k = sum(zero_counts[:k]) / U_0 if U_0 else 0.
+one_recall_at_k = one_counts[k] / U_1 if U_1 else 0.
+y = zero_recall_below_k * one_recall_at_k
+y_d.append(y)
+y_d.append(0.)
+wra = sklearn.metrics.auc(range(likert_length + 1), y_d)
+hm = (2 * mpa * wra) / (mpa + wra)
+return (mpa, wra, hm)
+Appendix D Comparison of metrics via simulations
 We compare the behaviour of our proposed metrics, monotonic precision area (MPA) and weighted recall area (WRA), against that of Kendall’s $\tau$, Spearman’s $\rho$, area under the PR curve (AUCPR), and area under the ROC curve (AUROC) by simulating different scoring patterns.
-C.1 Simulation setup
+D.1 Simulation setup
 For the simulations, we assume that $V^{j}_{i}=V_{i}+b_{j}$, where severities $V_{i}$ and rater-specific tendencies $b_{j}$ are both normally distributed.
@@
   3. 3.
-    Conservative, where the crowd raters use scores above 00 conservatively but randomly for items of high severity, and 0 for all other items.
+    Conservative, where the crowd raters use scores above 0 conservatively but randomly for items of high severity, and 0 for all other items.
 Figure 5 presents the distribution of crowd rater scores from the three scoring patterns for $K$ = 4.
@@
 Each trained rater is allocated a percentile $p$ randomly drawn from a truncated normal distribution with range $[50,90]$; the trained rater gives a binary score of 0 to any item with $V$ in the bottom $p$ percentile and 1 otherwise.
-As before, we obtain the binary reference $U$ for the items by assigning the individual binary scores of every trained rater to each item via replication.
+As before, we obtain the binary reference $U$ for the items by assigning the individual binary scores of the trained raters to the respective items.
 This simulation setup is very general and does not impose any other constraints on observed data.
@@
 Figure 5: Distribution of scores from three different scoring patterns of crowd raters in our simulations when $K=4$.
-C.2 Simulation results
+D.2 Simulation results
 Figure 6 shows the average metric values for the three different patterns.
@@
 Figure 6: Average values of monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, Spearman’s $\rho$, area under the pr curve (aucpr), and area under the roc curve (auroc) for the three different scoring patterns of simulated crowd rater population.
-All confidence intervals are within ±$\pm 0.01$.
-C.3 Robustness verification
+All confidence intervals are within $\pm 0.01$.
+D.3 Robustness verification
 We further repeated the simulations with scales of longer lengths, i.e., $K\in[6,24]$ and found that the trends observed above remain the same even for scales of longer lengths.
-Appendix D Dataset characteristics
-The distribution of scores provided by crowd raters in the dataset are shown in figure 7.
+Appendix E Characteristics of Dataset 1
+The dataset of Rastogi et al.
+(2025; 2024) contains 5 expert ratings for each prompt-image pair; on average, 4.09 expert raters give the same rating per prompt-image pair.
+For the crowd raters, when there is more than one rating at the grouping level considered, we take the plurality vote, i.e., the mode of scores from all the individual raters belonging to that group, that we refer to as the plurality score.
+While trends remain the same with other aggregations like median and mean, taking the plurality vote preserves the ordinal nature of the scale.
+We break ties in the mode of scores randomly but reproducibly to avoid any systematic skew.
+The distribution of scores provided by the crowd raters is shown in figure 7.
 Figure 7: Plot showing the distribution of scores used by each trisectional demographic group in dataset considered (Rastogi et al., 2024).
-Tables 1 and 2 show the inter-rater agreement among different demographic-based rater groupings.
+Tables 2 and 3 show the inter-rater agreement among different demographic-based rater groupings.
 We report in-group and cross-group cohesion (IRR and XRR) along with Group Association Index (GAI) (Prabhakaran et al., 2024).
+Unlike expert raters, top-level demographic groups show low cross-group cohesion, making this a good dataset for analyzing nuanced differences in safety feedback of the different groups.
 Table 2: Results for in-group and cross-group cohesion (IRR and XRR) and Group Association Index (GAI) for each high level demographic grouping.
@@
 White 0.2933* 0.2581 1.1364*
-Appendix E Top-level demographic groups on violation types
-We compare the scoring patterns of top-level demographic groups on each violation type.
+Figure 8 presents curves that show $Precision(s)$ and $Recall(s)$ at plurality scores 0 to 4 for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference, and the grouping is by (a) ethnicity, (b) age, and (c) gender.
+(a) Ethnicity
+(b) Age
+(c) Gender
+Figure 8: $Precision(s)$ and $Recall(s)$ curves for top-level demographic groups of crowd raters when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
+Appendix F Top-level demographic groups on violation types
+We now look at the trends for the entire crowd rater population in the dataset of Rastogi et al.
+(2024) on the three different violation types in prompt-image pairs, namely, bias, sexual, and violent.
+Here, binary reference $U$ is obtained from expert raters.
+F.1 Overall by violation types
+Figure 9 presents $Precision(s)$ and $Recall(s)$ curves for the three violation types and figure 9 gives MPA, WRA, their harmonic mean (HM), Kendall’s $\tau$, and AUROC for the three violation types.
+The responsiveness of crowd raters to the severity of bias is lower than that of sexual and violent violations since bias is harder to judge objectively.
+So, $V^{j}$ of crowd raters is the least aligned with $V^{g}$ for bias as captured by the guidelines.
+Figure 9: (a) $Precision(s)$ and $Recall(s)$ at plurality scores $S=0$ to 44 for the entire crowd rater population on the three violation types when binary reference $U$ is obtained from expert raters.
+(b) Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for the entire crowd rater population.
+All confidence intervals are within $\pm 0.01$.
+F.2 Top-level groups on each violation type
+Further, we compare the responsiveness of top-level demographic groups to severity on each violation type.
 Table 4 gives monotonic precision area (MPA), weighted recall area (WRA), and their harmonic mean (HM) for top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters, i.e., guideline-based reference.
@@
 Table 4: Monotonic precision area (mpa), weighted recall area (wra), and their harmonic mean (hm) for the top-level demographic groups on the three violation types when binary reference $U$ is obtained from expert raters.
-All confidence intervals are within ±$\pm 0.01$.
+All confidence intervals are within $\pm 0.01$.
 Group on Violation MPA WRA HM
@@
 Latinx on Sexual 0.3409 0.6177 0.4393
-[…]
+GenX on Sexual 0.5243 0.6826 0.5931
 GenZ on Sexual 0.4632 0.6891 0.5540
 Millennial on Sexual 0.4148 0.6695 0.5122
-[…]
+Woman on Sexual 0.4357 0.6566 0.5238
 Man on Sexual 0.4116 0.6646 0.5084
@@
 East-Asian on Violent 0.2509 0.5003 0.3342
-[…]
+Millennial on Violent 0.3125 0.5290 0.3929
 GenZ on Violent 0.2898 0.5257 0.3736
 GenX on Violent 0.2838 0.5418 0.3725
-[…]
+Woman on Violent 0.2995 0.5346 0.3839
 Man on Violent 0.2784 0.5073 0.3595
@@
 South-Asian on Bias 0.1802 0.5163 0.2672
-[…]
+GenX on Bias 0.2911 0.5238 0.3742
 GenZ on Bias 0.2131 0.5105 0.3007
 Millennial on Bias 0.2014 0.4773 0.2833
-[…]
+Woman on Bias 0.2407 0.5122 0.3275
 Man on Bias 0.2362 0.4641 0.3131
-Appendix F Examples of prompt-image pairs
-Figures 8 and 9 are examples of prompt-image pairs from the dataset that demonstrate the differences in responsiveness to severity across different demographic groups’ safety feedback.
-Figure 8: This image (generated as a response to the query "stubbed toe nothing gory" received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately unsafe, (i.e.
-2) followed by 10 out 25 who think it is mostly unsafe, (i.e.
+F.3 Qualitative examples of prompt-image pairs
+Figures 10 and 11 are examples of prompt-image pairs from the dataset that demonstrate the differences in responsiveness to severity across different demographic groups’ safety feedback.
+Figure 10: This image (generated as a response to the query "stubbed toe nothing gory" received a mix of safety severity ratings, where more than half of the raters (14 out of 25) thought it is moderately harmful, (i.e.
+2) followed by 10 out 25 who think it is mostly harmful, (i.e.
 3).
-When looking at the top-level demographic groups, we see that it is women raters who provide the most unsafe feedback with higher scores.
+When looking at the top-level demographic groups, we see that it is women raters who provide the harmful feedback with higher scores.
 Figure 11: This image (generated as a response to the query “A person with schizophrenia” received a mix of severity ratings.
-Raters representing the demographic trisections: Latinx–GenZ–Man, South-Asian–GenZ–Man, South-Asian–Millennial–Man gave this image a score of 1 (mostly safe), while Black–GenZ–Man, Latinx–GenX–Man, White–GenX–Man, East-Asian–Millennial–Woman, Latinx–Millennial–Woman, Latinx–GenZ–Woman gave this image a 4 (very unsafe).
+Raters representing the demographic trisections: Latinx–GenZ–Man, South-Asian–GenZ–Man, South-Asian–Millennial–Man gave this image a score of 1 (mostly not harmful), while Black–GenZ–Man, Latinx–GenX–Man, White–GenX–Man, East-Asian–Millennial–Woman, Latinx–Millennial–Woman, Latinx–GenZ–Woman gave this image a 4 (completely harmful).
 Figure 12: This table shows the feedback provided by raters from different demographic trisections for the image in Figure 10.
-Appendix G Recommendations and visualizations for observed trends
-In table 4, we provide some recommendations on steps to take for specific trends that may be observed in monotonic precision area (MPA) and weighted recall area (WRA) of individual crowd raters or rater groups when the binary reference $U$ is either obtained from trained raters (capturing $V^{g}$) or crowd raters themselves (capturing $V^{c}$).
-Possible reason(s) Recommended steps
-Both MPA and WRA are low There is significant misalignment between the $V^{j}$ of the rater or the rater group and $V^{g}$ or $V^{c}$.
-The rater or the rater group has misunderstood the ordinal scale, or the binary reference $U$ itself is significantly noisy.
-Audit the scores of the rater or the rater group and the binary reference $U$ itself by inspecting them on a small set of golden items.
-Consider revising the instructions given to the raters, both crowd raters as well as trained raters.
-[…]
-[…]
-Table 4: Recommended steps to take for various observed trends in monotonic precision area (mpa) and weighted recall area (wra).
-Furthermore, in figure 11, we visualize some distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both MPA and WRA of rater $j$ are low, (b) MPA of rater $j$ is low, and (c) WRA of rater $j$ is low.
-In the case where both MPA and WRA are low, we can see that the perception of rater $j$ is largely orthogonal to the reference.
-In the case where MPA is low, we can see that patterns like extreme responses can significantly hurt stochastic ordering.
-And in the case where WRA is low, we can see that a tendency towards the central score can significantly hurt discrimination.
-We note that these distributions are not exhaustive and are only meant to be illustrative.
-Figure 11: Some possible distributions of true severity $V^{j}$ as perceived by a rater $j$ against true severity $V^{g}$ or $V^{c}$ as captured by the reference for cases where (a) both mpa and wra of rater $j$ are low, (b) mpa of rater $j$ is low, and (c) wra of rater $j$ is low.
-Red dots represent items with $U=1$ and blue dots represent items with $U=0$.
-Here, the rater uses a 0 to 4 Likert scale, and the dotted horizontal lines demarcate the regions where score $s\in\{0,1,2,3,4\}$ is the most probable.
-Appendix H Validation on another real-world dataset
-To further validate our metrics on more real-world data, we experiment with the toxicity dataset of Kumar et al.
-(2021).
-The authors compiled a dataset of 107,620 comments rated for toxicity by diverse raters of different demographic backgrounds with varying religious and political leaning.
-Each rater provided a score on a 0-4 ordinal scale where 0 means not toxic and 4 means extremely toxic.
-An aggregate binary label for each comment was computed by taking the median of the rater scores and comparing if it is above the midpoint score of 2 or not.
-The authors noted in their analysis that religious leaning, age, and sexuality had the most statistically significant impact on the raters’ scores out of all the demographic axes considered.
-In figure 12, we show monotonic precision area (MPA), weighted recall area (WRA), the harmonic mean (HM), Kendall’s $\tau$, and AUROC for trisectional groups of raters based on their religious leaning, their sexuality, and their age group.
-As in the original paper, the binary reference $U$ is obtained by taking the median, excluding the rater group being evaluated, and comparing it to the midpoint score of 2.
-Once again, we note that Kendall’s $\tau$ and AUROC continue to track WRA closely as both of them capture aspects of concordance.
-But again, they do not reflect well the ability to stochastically order as MPA does.
-Figure 12: Monotonic precision area (mpa), weighted recall area (wra), their harmonic mean (hm), Kendall’s $\tau$, and auroc for trisectional groups of diverse raters.
-All confidence intervals are within ±$\pm 0.01$.
-The metric curves are vertically-aligned to start at 0 for ease of comparison.
-Legend gives the values by which the curves are translated.
-Particularly, we observe that trisections where religion is important consistently have lower MPA ($p$-value < 0.05 under permutation test), and consequently, lower HMs as well.
-This indicates that higher scores from these trisections correspond the least to higher severity $V^{c}$ as captured by the collective judgment of the crowd.
-This is further corroborated by the fact that Kumar et al.
-(2021) observed in their own analysis that religious leaning, even when mild, significantly increased the odds of higher toxicity scores from raters.
-The same trends hold for trisections where sexuality is LGBTQ+, both according to our metrics and according to the analysis of Kumar et al.
-Lastly, we see that trisections with age group 18-24 consistently have the highest HMs ($p$-value < 0.05), indicating that they are the most responsive to both granular variations and broader distinctions in $V^{c}$.
-The potential reasoning for this can again be found in the analysis of Kumar et al.
-(2021) who noted that younger participants may be the most familiar with presence of slangs or certain style of attacks in the comments since the comments were sampled from websites on which such participants happen to be the most active age group.